### v1

# Accuracy of Discretely Sampled Stochastic Policies in Continuous-time Reinforcement Learning

Abstract.
Stochastic policies are widely used in continuous-time reinforcement learning algorithms. However, executing a stochastic policy and evaluating its performance in a continuous-time environment remain open challenges. This work introduces and rigorously analyzes a policy execution framework that samples actions from a stochastic policy at discrete time points and implements them as piecewise constant controls. We prove that as the sampling mesh size tends to zero, the controlled state process converges weakly to the dynamics with coefficients aggregated according to the stochastic policy. We explicitly quantify the convergence rate based on the regularity of the coefficients and establish an optimal first-order convergence rate for sufficiently regular coefficients. Additionally, we show that the same convergence rates hold with high probability concerning the sampling noise, and further establish a $1/2$-order almost sure convergence when the volatility is not controlled. Building on these results, we analyze the bias and variance of various policy evaluation and policy gradient estimators based on discrete-time observations. Our results provide theoretical justification for the exploratory stochastic control framework in [H. Wang, T. Zariphopoulou, and X.Y. Zhou, J. Mach. Learn. Res., 21 (2020), pp. 1-34].
Key words and phrases:
policy execution, stochastic policy, continuous-time reinforcement learning, exploratory control, piecewise constant control, convergence rate
1. Introduction
Continuous-time RL
Reinforcement learning (RL) has catalyzed significant changes across various domains by attacking complex sequential decision-making challenges in a data-driven way. The theoretical framework for analyzing RL problems and designing RL algorithms has been predominantly based on the discrete-time Markov decision processes [5, 30]. However, many practical applications require real-time interaction with environments where time progresses continuously, such as autonomous driving, stock trading, and vehicle routing. Recent studies [33, 38, 20, 12] have shown that certain algorithms originally designed for discrete-time RL can become unstable when applied in high-frequency settings. This highlights the need for a new theoretical framework for designing RL algorithms that are robust to time discretization.
To address this issue, many suggest formulating the problem in a continuous-time framework from the outset and only introducing time discretization at the final implementation stage. Here, time discretization refers to the frequency at which data is collected and actions are updated, while the underlying system itself evolves continuously according to its intrinsic dynamics. Early studies on continuous-time RL have primarily focused on deterministic systems, where state variables follow ordinary differential equations; see, e.g., [1, 9, 33, 26, 22].
Exploratory control and stochastic policy
Recently, [36] proposed a novel exploratory control formulation with stochastic policies for continuous-time RL problems, which has served as the foundation for designing various RL algorithms; see, e.g., [18, 19, 20, 15, 8, 11, 37, 12, 29, 32].
This control formulation in [36] has three key features: (1) The agent incorporates stochastic policies into the control problem, which models exploration in RL. These policies are mappings from the state space to distributions over the action space, aligning with the notion of “relaxed control” in classical control theory. (2) The agent continuously samples actions from a stochastic policy, and the state process evolves according to a new diffusion process, with its coefficients integrated with respect to the sampling distribution specified by the stochastic policy. This auxiliary state dynamic is often referred to as the exploratory or aggregated dynamics. (3) The optimal policy remains stochastic due to the inclusion of an additional entropy regularizer in the cost objective, in contrast to the deterministic/strict policies typically found in standard control problems.
Execution of stochastic policies
However, despite its success in algorithm design, the use of stochastic policies in continuous-time RL problems introduces theoretical and practical gaps. Specifically, the connection between exploratory dynamics with aggregated coefficients and state dynamics controlled by frequent sampling from a stochastic policy remains unjustified. In fact, interacting with a diffusion process by continuously sampling from a stochastic policy requires generating uncountably many independent random variables, which raises measure-theoretical issues and renders the resulting controlled state dynamics ill-defined (see the end of Section 2 for more details).
To address this issue, [32] proposes approximating exploratory dynamics by sampling actions from a stochastic policy only at discrete time points. This approach naturally aligns with the practical implementation of stochastic policies. However, for algorithms designed based on aggregated dynamics, executing a stochastic policy in discrete time introduces a bias that depends on the sampling frequency. This leads to the following natural question:
What is the precise impact of sampling frequency on the performance of a stochastic policy and the corresponding learning algorithms?
To the best of our knowledge, this question has only been addressed in the context of linear-quadratic (LQ) RL problems with Gaussian policies [12, 32]. In this special setting, due to the LQ structure, the analysis is simplified by examining the discrepancy between the first and second moments of the aggregated dynamics and the controlled dynamics with sampled actions. Unfortunately, these techniques developed specifically for LQ RL problems are clearly not applicable to general RL problems. In particular, it is essential to quantify the difference between the entire distributions of two state processes.
Our contributions
This paper analyzes the accuracy of the discrete-time execution of stochastic policies for general diffusion processes and quantifies its impact on the associated learning algorithms.
  • We show that when actions are sampled from a stochastic policy at a discrete time grid $\mathscr{G}$ and executed as a piecewise constant control process, the resulting state process, referred to as the sampled dynamics, converges weakly to the aggregated dynamics as the time stepsize approaches zero. In particular, under sufficient regular coefficients, we show that the convergence rate is of first order $O(|\mathscr{G}|)$, with $|\mathscr{G}|$ being the time stepsize (Theorem 4.1). This matches the optimal convergence rate for standard piecewise constant control approximations [16]. We further quantify the precise weak convergence rate for cases with less regular coefficients (Theorem 4.2).
    The convergence results rigorously justify the exploratory control framework proposed in [36] (see Section 2). The weak convergence rate will be used to analyze the performance of learning algorithms based on observations from the sampled dynamics.
  • We prove that the sampled dynamics, in fact, converges almost surely with respect to the randomness in policy execution. Specifically, we show that the weak convergence order to the aggregated dynamics holds with high probability with respect to the randomness used to sample actions (Theorem 4.3). We further show that when the volatility is uncontrolled, the sampled dynamics converges almost surely to the aggregated dynamics with an order of $\mathcal{O}(|\mathscr{G}|^{1/2})$, for each fixed realization of the Brownian motion (Theorem 4.4). The analysis leverages concentration inequalities for bounded and sub-exponential martingale difference sequences to control the concentration behavior of the sampling errors in terms of the sampling frequency (see the discussions below Theorems 4.3 and 4.4).
    The almost sure convergence result indicates that the convergence of the sampled dynamics to the aggregated dynamics can be interpreted as a law of large numbers in time, i.e., the convergence can be achieved by refining the sampling frequency while fixing the realization of the system noise. Note that such an almost sure convergence fails when the volatility is controlled, as shown in Example 4.6.
  • We apply the weak convergence rate to quantify the error incurred when implementing continuous-time RL algorithms based on observations from sampled dynamics. In particular, we show that evaluating the reward involving aggregated dynamics by discrete sampling and piecewise constant execution has a bias of the order $O(|\mathscr{G}|)$. Calculating the martingale orthogonality conditions proposed in [18] using discretely sampled stochastic policies also introduces a bias of the same order, $O(|\mathscr{G}|)$. The policy gradient representation developed in [19] is an informal integration that involves the realized actions and the state variables. We rigorously show that using the forward Euler scheme to calculate such terms with discretely sampled stochastic policies is a biased estimator for the policy gradient, and the bias is $O(|\mathscr{G}|)$ and variance is bounded. The q-learning with quadratic variation penalty for the risk-sensitive RL proposed in [17] is also shown to have a similar order of discretization error.
Our approaches and most related works
This paper quantifies the accuracy of the discrete-time execution of a stochastic policy and its impact on continuous-time RL algorithms. The accuracy of policy execution has been studied in [14, 32] for LQ problems with Gaussian policies. However, as mentioned earlier, the techniques developed in [14, 32] are not applicable to RL problems with general state dynamics and stochastic policies considered in this paper.
For general state dynamics, the weak convergence of sampled dynamics has recently been studied by [4, 3]. They adopt a different definition of sampled dynamics, treating it as being driven by a random measure, and establish the weak convergence of the state process without quantifying the convergence order. In contrast to the time-varying control considered in [4, 3], we analyze sampled dynamics controlled by a piecewise constant control process. Moreover, we take a completely different approach and establish the weak convergence rate under various regularity conditions by leveraging the solution regularity of the associated partial differential equations (PDEs). Beyond weak convergence, we also establish convergence in the almost sure sense with respect to the policy execution noise. These strong convergence results require a careful analysis of the concentration properties of the state dynamics concerning the policy execution noise.
In addition to discrete sampling error, we also analyze the error incurred when approximating continuous-time RL algorithms using discrete-time observations. [18] analyzes a similar error but assumes that observations are generated from the aggregated dynamics. [2] analyzes both errors for LQ problems with uncontrolled volatility.
Our analysis of the policy execution error is fundamentally different from classical numerical analysis of time discretizations for stochastic differential equations (SDEs). In particular, numerical analysis of SDEs focuses on the error caused by discretizing the system noise, such as Brownian motion, while both the SDE and its time-discretized version have the same coefficients. In our setting, we treat the controlled SDE as a black-box environment evolving continuously and focus on the error introduced by the piecewise constant execution of a stochastic policy. As a result, the environment noise is not discretized, but the limiting SDE (the aggregated dynamics) and the sampled dynamics have different coefficients. The former involves aggregated coefficients, while the latter has random coefficients determined by the specific procedure used to sample the policy.
The rest of the paper is organized as follows. Section 2 briefly reviews the use of stochastic policies in continuous-time RL. Section 3 presents the problem setup and introduces the sampled dynamics rigorously. Section 4 presents the main convergence results of the sampled dynamics, and Section 5 applies them to analyze continuous-time RL algorithms. Finally, we conclude in Section 6. The proofs of all statements are included in Section 7.
2. Preliminary
In this section, we review the exploratory control formulation and the use of stochastic policies for continuous-time RL algorithms. The precise assumptions and setup can be found in Section 3.
2.1. Classical stochastic control
Recall that in a classical stochastic control problem, an agent aims to control the state process and optimize an objective functional through an adapted control process taking values in a Polish space $A$. Specifically, given an admissible control process $\{a_{t}\}_{t\geq 0}$, the corresponding state process $\{X_{t}\}_{t\geq 0}$ is governed by the following dynamics: [^1] To simplify the presentation, we assume that the state processes have a deterministic initial condition. However, our analysis and results extend naturally to cases with a random initial condition, provided it is independent of the Brownian motion $W$ and satisfies appropriate integrability conditions. We also assume that the Brownian motion’s dimension equals the state variable’s dimension; otherwise, we can introduce more Brownian motions or state variables to make them equal.
$\mathrm{d}X_{t}=b(t,X_{t},a_{t})\mathrm{d}t+\sigma(t,X_{t},a_{t})\mathrm{d}W_{t},\quad t\in[0,T];\quad X_{0}=x_{0},$ (2.1)
where $x_{0}\in{\mathbb{R}}^{d}$ is a given initial state, $b:[0,T]\times\mathbb{R}^{d}\times A\to\mathbb{R}^{d}$ and $\sigma:[0,T]\times\mathbb{R}^{d}\times A\to\mathbb{R}^{d\times d}$ are given drift and volatility coefficients, respectively, and $W$ is an $d$-dimensional Brownian motion supported on a probability space $(\Omega,\mathcal{F},\mathbb{P})$. The agent’s objective is to maximize the following objective
$\mathbb{E}\left[\int_{0}^{T}r(t,X_{t},a_{t})\mathrm{d}t+h(X_{T})\right]$ (2.2)
over all admissible controls $\{a_{t}\}_{t\geq 0}$, where $r:[0,T]\times{\mathbb{R}}^{d}\times A\to{\mathbb{R}}$ is a given instantaneous reward function, and $h:{\mathbb{R}}^{d}\to{\mathbb{R}}$ is a given terminal reward function.
It is known that under suitable conditions (see [39]), an optimal control for (2.2) exists and is given by a feedback form, i.e., there exists a measurable function $\bm{u}:[0,T]\times\mathbb{R}^{d}\to A$ such that the optimal control is obtained by $a_{t}=\bm{u}(t,X_{t}^{\bm{u}})$ for all $t\in[0,T]$, where $\{X_{t}^{\bm{u}}\}_{t\geq 0}$ satisfies the following dynamics:
$\mathrm{d}X_{t}=b(t,X_{t},\bm{u}(t,X_{t}))\mathrm{d}t+\sigma(t,X_{t},\bm{u}(t,X_{t}))\mathrm{d}W_{t},\quad t\in[0,T];\quad X_{0}=x_{0},$ (2.3)
The function can be determined using the coefficients $b,\sigma,r,$ and $h$. Once determined, the optimal control can be implemented by evaluating ${\bm{u}}$ at the current state of the system.
2.2. RL, stochastic policy and exploratory control
In the RL problem, the agent seeks to optimize the objective (2.2) by directly interacting with the dynamics (2.1), experimenting with different actions, and refining their strategies based on the observed responses. Unlike traditional control settings, the agent treats (2.1) as a black-box environment evolving in continuous time and learns an optimal policy solely from observations, without direct access to the coefficients $b,\sigma,r,$ and $h$.
To avoid getting trapped in local optima, a key feature of RL is explicitly exploring the environment through randomized actions. A common approach is to use stochastic policies $\bm{\pi}:[0,T]\times{\mathbb{R}}^{d}\to\mathcal{P}(A)$, which maps the current state to a probability distribution over the action space. Given a stochastic policy $\bm{\pi}$, the agent interacts with the environment by sampling noisy actions from $\bm{\pi}$ and observing the resulting system states. Over time, the agent refines the stochastic policy to optimize the objective.
To describe the performance of a stochastic policy $\bm{\pi}$, [36] suggests that the associated state process $\tilde{X}$ evolves according to the following dynamics (called the exploratory dynamics therein):
${\mathrm{d}}X_{t}=\tilde{b}^{\bm{\pi}}(t,X_{t})\mathrm{d}t+\tilde{\sigma}^{\bm{\pi}}(t,X_{t})\mathrm{d}W_{t},\quad t\in[0,T];\quad X_{0}=x_{0},$ (2.4)
with the coefficients $\tilde{b}^{\bm{\pi}}:[0,T]\times{\mathbb{R}}^{d}\to{\mathbb{R}}^{d}$ and $\tilde{\sigma}^{\bm{\pi}}:[0,T]\times{\mathbb{R}}^{d}\to{\mathbb{R}}^{d\times d}$ given by
$\tilde{b}^{\bm{\pi}}(t,x)\coloneqq\int_{A}b(t,x,a)\bm{\pi}(\mathrm{d}a|t,x),\quad\tilde{\sigma}^{\bm{\pi}}(t,x)\coloneqq\sqrt{\int_{A}(\sigma\sigma^{\top})(t,x,a)\bm{\pi}({\mathrm{d}}a|t,x)},$ (2.5)
where $(\cdot)^{1/2}$ is the matrix square root of positive semidefinite matrices, and $(\cdot)^{\top}$ is the matrix transpose. Moreover, the associated objective function is given by:
$\mathbb{E}^{W}\left[\int_{0}^{T}\tilde{r}(t,\widetilde{X}_{t})\mathrm{d}t+h(\widetilde{X}_{T})\right],\quad\textnormal{with $\tilde{r}(t,x)\coloneqq\int_{A}r(t,x,a)\bm{\pi}(\mathrm{d}a|t,x)$. }$ (2.6)
The dynamics (2.4) and the objective (2.6) are derived heuristically in [36] under the assumption that the agent interacts with the system (2.1) by continuously sampling independent random actions from the policy $\bm{\pi}$. As a result of these random actions, a “law-of-large-numbers” heuristic suggests that the infinitesimal change in the state at time $t$ is determined by the coefficients in (2.5), aggregated over the sampling distribution defined by $\bm{\pi}$. Likewise, the instantaneous reward is also aggregated with respect to this sampling distribution. In the sequel, we refer to (2.4) as the aggregated dynamics to highlight this aggregating effect.
However, the rigorous justification of the connection between (2.4) and the state dynamics resulting from the adoption of a stochastic policy $\bm{\pi}$ in (2.1), as well as the relationship between (2.6) and the original reward functional (2.2), remains to be established. Indeed, interacting with the SDE (2.1) by continuously sampling from a stochastic policy creates measure-theoretical issues. As already pointed out in [32, Remark 2.1], it is impossible to construct a family of non-constant random variables $(\xi_{t})_{t\in[0,1]}$ such that $(\xi_{t})_{t\in[0,1]}$ is (essentially) pairwise independent and $t\mapsto\xi_{t}$ is Lebesgue measurable. This implies that if one controls (2.1) by continuously generating independent actions, the resulting coefficients are not progressively measurable, rendering the conventional stochastic integral ill-defined; see [3] for more discussion. Moreover, it is practically infeasible to generate these independent samples continuously.
Therefore, both from theoretical and practical perspectives, evaluating the performance of a stochastic policy $\bm{\pi}$ requires discretely sampling actions from the policy and applying them to (2.1), as suggested by [32]. This work quantifies the accuracy of such a discrete execution of a stochastic policy and its impact on the corresponding RL algorithms.
3. Probabilistic Setup for Aggregated and Sampled Dynamics
This section introduces a rigorous probabilistic framework for the aggregated dynamics under a given stochastic policy and the associated sampled dynamics with a finite number of intervention times, as outlined in Section 2.
3.1. Aggregated dynamics
Fix a probability space $(\Omega^{W},\mathcal{F}^{W},\mathbb{P}^{W})$ on which a $d$-dimensional Brownian motion $\{W_{t}\}_{t\geq 0}$ is defined, and let ${\mathbb{F}}^{W}\coloneqq\{\mathcal{F}_{t}^{W}\}_{t\geq 0}$ be the $\mathbb{P}^{W}$-completion of the filtration generated by $W$, which satisfies the usual conditions of completeness and right-continuity. Throughout this paper, we impose the following regularity condition on the action space $A$, the drift and diffusion coefficients $b$ and $\sigma$, and the stochastic policy $\bm{\pi}$.
H. 1.
  1. (1)
    $({A},d_{A})$ is a Polish space equipped with the Borel $\sigma$-algebra $\mathcal{B}({A})$, and $a_{0}\in A$;
  2. (2)
    $b:[0,T]\times\mathbb{R}^{d}\times{A}\to\mathbb{R}^{d}$ and $\sigma:[0,T]\times\mathbb{R}^{d}\times{A}\to\mathbb{R}^{d\times d}$ are measurable functions. There exists $C\geq 0$ such that for all $\varphi\in\{b,\sigma\}$, $t\in[0,T]$, $x,x^{\prime}\in\mathbb{R}^{d}$ and $a\in{A}$,
    $|\varphi(t,x,a)-\varphi(t,x^{\prime},a)|\leq C|x-x^{\prime}|,\quad|\varphi(t,x,a)|$ $\leq C(1+|x|+d_{A}(a,a_{0}));$
  3. (3)
    $\bm{\pi}:[0,T]\times\mathbb{R}^{d}\times\mathcal{B}({A})\to[0,1]$ is a probability kernel. For all $p\geq 2$, there exists $C_{p}\geq 0$ such that for all $(t,x)\in[0,T]\times{\mathbb{R}}^{d}$, $\int_{A}d_{A}(a,a_{0})^{p}\bm{\pi}({\mathrm{d}}a|t,x)\leq C_{p}(1+|x|^{p})$.
Condition (H.12) imposes the standard Lipschitz regularity on the system coefficients commonly used in the stochastic control literature (see, e.g., [10, 39]), and Condition (H.13) requires mild growth and integrability properties for the stochastic policy. Under (H.1), $\tilde{b}^{\bm{\pi}}$ and $\tilde{\sigma}^{\bm{\pi}}$ in (2.5) are well-defined measurable functions and satisfy $|\tilde{b}^{\bm{\pi}}(t,x)|+|\tilde{\sigma}^{\bm{\pi}}(t,x)|\leq C(1+|x|)$ for all $(t,x)\in[0,T]\times{\mathbb{R}}^{d}$.
We assume that $\tilde{b}^{\bm{\pi}}$ and $\tilde{\sigma}^{\bm{\pi}}$ are sufficiently regular so that the aggregated dynamic (2.4) is well-posed.
H. 2.
The functions $\tilde{b}^{\bm{\pi}}$ and $\tilde{\sigma}^{\bm{\pi}}$ are sufficiently regular so that (2.4) admits a unique strong solution $\tilde{X}$ on the filtered probability space $(\Omega^{W},\mathcal{F}^{W},{\mathbb{F}}^{W},\mathbb{P}^{W})$.
Here, we focus on strong solutions of (2.4) rather than weak solutions, which allows us to identify $W$ as the only driving noise in the state process $\tilde{X}$. This simplifies the construction of the sampled dynamics and the subsequent error estimates in terms of the sampling frequency. Under (H.1), Condition (H.2) holds if $\tilde{b}^{\bm{\pi}}$ and $\tilde{\sigma}^{\bm{\pi}}$ are locally Lipschitz continuous in the state variable (cf. [27, Theorem 3.4]) and can be ensured via suitable regularity conditions on $\bm{\pi}$ (e.g., [19, Lemma 2]). Some typical sufficient conditions on $b,\sigma$ and $\bm{\pi}$ to guarantee these regularity conditions on $\tilde{b}^{\bm{\pi}}$ and $\tilde{\sigma}^{\bm{\pi}}$ are given in Examples 4.1, 4.2 and 4.3.
3.2. Sampled dynamics with randomized actions
We proceed to define the state process with random actions sampled according to the stochastic policy $\bm{\pi}$. Rather than focusing on a specific sampling procedure for $\bm{\pi}$, we adopt a more abstract formulation of the sampling process defined as follows.
Definition 3.1.
Assume the notation of (H.1). We say a tuple $(\Omega^{\xi},\mathcal{F}^{\xi},\mathbb{P}^{\xi},E,\xi,\phi)$ a sampling procedure of the policy $\bm{\pi}$ if $(\Omega^{\xi},\mathcal{F}^{\xi},\mathbb{P}^{\xi})$ is a complete probability space, $(E,\mathcal{B}(E))$ is a Borel space, $\xi:\Omega^{\xi}\to E$ is a random variable and $\phi:[0,T]\times{\mathbb{R}}^{d}\times E\to A$ is a measurable function such that for all $(t,x)\in[0,T]\times{\mathbb{R}}^{d}$, $\Omega^{\xi}:\omega\mapsto\phi(t,x,\xi(\omega))\in A$ has the distribution $\bm{\pi}({\mathrm{d}}a|t,x)$ under the measure $\mathbb{P}^{\xi}$.
By Definition 3.1, $(E,\xi,\phi)$ provides a framework for executing the policy $\bm{\pi}$ by sampling a random action $a_{t}\coloneqq\phi(t,X_{t},\xi)$ from the distribution $\bm{\pi}({\mathrm{d}}a|t,X_{t})$ at a given time $t\in[0,T]$ and state $X_{t}$. This action, conditioned on the $\sigma$-algebra $\sigma\left\{(X_{s})_{s\in[0,t]}\right\}$, is independent of both $(X_{s})_{s\in[0,t]}$ and the underlying Brownian motion $W$.
Remark 3.1.
As $A$ is a Polish space, a sampling procedure of a probability kernel $\bm{\pi}$ always exists. Indeed, by [21, Lemma 2.22], it suffices to take $E=[0,1]$ and $\xi$ as a uniform random variable on $E$. This approach has been utilized in [11, 4, 3]. However, the corresponding function $\phi$ is guaranteed to exist only through a measure-theoretic argument, making it difficult to implement such a sampling procedure.
Here we work with a general Borel space $E$, instead of restricting $E$ to $[0,1]$. This generalization accommodates commonly used sampling procedures that generate actions using multiple independent normal random variables, as seen in applications such as multivariate Gaussian policies [32] or suitable Markov chain Monte Carlo methods. In Section 4, we establish the weak convergence of the control-randomized SDE under the mere measurability assumption of $\phi$. Additional regularity of the function $\phi$ will be imposed to obtain strong convergence results.
Now, fix a sampling procedure $(\Omega^{\xi},\mathcal{F}^{\xi},\mathbb{P}^{\xi},E,\xi,\phi)$ of $\bm{\pi}$, let ${\mathbb{N}}_{0}={\mathbb{N}}\cup\{0\}$ and let $(\Omega^{\xi_{n}},\mathcal{F}^{\xi_{n}},\mathbb{P}^{\xi_{n}},\xi_{n})_{n\in{\mathbb{N}}_{0}}$ be independent copies of $(\Omega^{\xi},\mathcal{F}^{\xi},\mathbb{P}^{\xi},\xi)$. Consider a probability space of the following form:
$(\Omega,\mathcal{F},\mathbb{P})\coloneqq\bigg{(}\Omega^{W}\times\prod_{n=0}^{\infty}\Omega^{\xi_{n}},\mathcal{F}^{W}\otimes\bigotimes_{n=0}^{\infty}\mathcal{F}^{\xi_{n}},\mathbb{P}^{W}\otimes\bigotimes_{n=0}^{\infty}\mathbb{P}^{\xi_{n}}\bigg{)},$ (3.1)
where $(\Omega^{W},\mathcal{F}^{W},\mathbb{P}^{W})$ supports the Brownian motion $W$ in (2.4), and for each $n\in{\mathbb{N}}_{0}$, $(\Omega^{\xi_{n}},\mathcal{F}^{\xi_{n}},\mathbb{P}^{\xi_{n}})$ supports the random variable $\xi_{n}$ used to generate the random control at the grid point $t_{n}$. With a slight abuse of notation, we complete the probability space $(\Omega,\mathcal{F},\mathbb{P})$ by augmenting $\mathcal{F}$ with $\mathbb{P}$-null sets, and extend canonically any random variable on a marginal space to the whole space $\Omega$, e.g., $W(\omega^{W},\omega^{\xi})=W(\omega^{W})$ and $\widetilde{X}(\omega^{W},\omega^{\xi})=\widetilde{X}(\omega^{W})$ for all $\omega^{W}\in\Omega^{W}$ and $\omega^{\xi}\in\prod_{n=0}^{\infty}\Omega^{\xi_{n}}$.
Given a time grid $\mathscr{G}=\{0=t_{0}<\ldots<t_{n}=T\}$ of $[0,T]$, we consider interacting with the state dynamics (2.1) by sampling actions at the grid points in $\mathscr{G}$ according to the policy $\bm{\pi}$. More precisely, we consider the dynamics such that for all $i=0,\ldots,n-1$ and all $t\in[t_{i},t_{i+1}]$,
$\begin{split}X_{t}&=X_{t_{i}}+\int_{t_{i}}^{t}b(s,X_{s},a_{t_{i}}){\mathrm{d}}s+\int_{t_{i}}^{t}\sigma(s,X_{s},a_{t_{i}}){\mathrm{d}}W_{s},\quad\textnormal{with $a_{t_{i}}=\phi(t_{i},X_{t_{i}},\xi_{i})$};\quad X_{0}=x_{0},\end{split}$ (3.2)
which will be referred to as the sampled dynamics [^2] This definition slightly differs from the definition of the “grid sampling SDE” in [3, Section 3] even though both rely on drawing finitely many samples on the grid point. In our definition, the action sequence $a_{t}$ is a piecewise constant and is given by a sampling function at the grid points $a_{t_{i}}=\phi(t_{i},X_{t_{i}},\xi_{i})$. By contrast, in the grid sampling SDE in [3], the action sequence is evolving as $a_{t}=\phi(t,X_{t},\xi_{t_{i}})$, where the random source $\xi_{t_{i}}$ is piecewise constant. in the subsequent analysis. For notational convenience, we write (3.2) in the following equivalent form:
$\mathrm{d}X_{t}=b(t,X_{t},a_{\delta(t)})\mathrm{d}t+\sigma(t,X_{t},a_{\delta(t)})\mathrm{d}W_{t},\quad t\in[0,T];\quad X_{0}=x_{0}$ (3.3)
with $\delta(t)\coloneqq t_{i}$ for $t\in[t_{i},t_{i+1})$, and $a_{t_{i}}$ given in (3.2). The dynamics (3.3) can be viewed as an SDE with random coefficients.
Remark 3.2.
Although the definition of the sampled dynamics depends on the sampling scheme $\phi$ and $\xi$, the weak convergence rate of the sampled dynamics to the aggregated dynamics depends only on the properties of the stochastic policy $\bm{\pi}$, rather than the specific choice of $\phi$ and $\xi$ (Sections 4.1 and 4.2). The properties of $\phi$ will be used to quantify the pathwise error between the sampled dynamics and the aggregated dynamics (Section 4.3).
Note that the sampled dynamics evolve continuously over time while the control process remains constant within each subinterval. In particular, a random action $a_{t_{i}}$ is generated at $t_{i}$ and applied to the system over the interval $[t_{i},t_{i+1})$ before being updated to the next action $a_{t_{i+1}}$. This aligns with the RL setting, where the agent observes a continuous-time system at discrete time points and adjusts their actions based on these observations.
The following lemma shows that the sampled dynamics (3.3) admits a unique strong solution $X^{\mathscr{G}}$ which is adapted to the filtration generated by both the Brownian motion $W$ and the execution noise $\xi$. The proof is given in Appendix A.
Lemma 3.1.
Suppose (H.1) holds. Let $x_{0}\in{\mathbb{R}}^{d}$, $(\Omega,\mathcal{F},{\mathbb{F}},\mathbb{P})$ be the probability space defined in (3.1), and ${\mathbb{F}}=(\mathcal{F}_{t})_{t\geq 0}$ be the filtration such that $\mathcal{F}_{t}\coloneqq\sigma\{(W_{s})_{s\leq t},(\xi_{i})_{i=0}^{\infty}\}\vee\mathcal{N}$, where $\mathcal{N}$ be the $\mathbb{P}$-null set. Then for any grid $\mathscr{G}$, (3.3) admits a unique ${\mathbb{F}}$-adapted strong solution $X^{\mathscr{G}}$ on the space $(\Omega,\mathcal{F},\mathbb{P})$. Moreover, for all $p\geq 2$, there exists a constant $C_{p}\geq 0$ such that for all grids $\mathscr{G}$ and all $t\in[0,T]$, $\mathbb{E}\left[|X^{\mathscr{G}}_{t}|^{p}\right]\leq C_{p}(1+|x_{0}|^{p})e^{C_{p}t}$, and $X^{\mathscr{G}}_{t}$ is adapted to $\mathcal{G}_{t}\coloneqq\sigma\{(W_{s})_{s\leq t},(\xi_{i})_{t_{i}<t}\}\vee\mathcal{N}$.
4. Error Estimates for the Sampled Dynamics
In this section, we analyze the convergence rate of the sampled dynamics (3.3) to the aggregated dynamics (2.4) under various criteria as the time grid $\mathscr{G}$ is refined. In particular, we quantify the weak error for quantities involving expectations over the system noise $W$ (but not necessarily over the execution noises $(\xi_{i})_{i=0}^{\infty}$), and the pathwise error $\tilde{X}-X^{\mathscr{G}}$. To this end, we define the mesh size of a time grid $\mathscr{G}=\{0=t_{0}<\ldots<t_{n}=T\}$ by $|\mathscr{G}|=\max_{0\leq i\leq n-1}(t_{i+1}-t_{i})$.
4.1. Weak error with expectations over execution noise
We start by analyzing the weak convergence rate of $X^{\mathscr{G}}$ to $\tilde{X}$ under the measure $\mathbb{P}$. That is, for a suitable test function $f:{\mathbb{R}}^{d}\to{\mathbb{R}}$, we quantify the error $\mathbb{E}[f(\widetilde{X}_{s})]-\mathbb{E}[f(X^{\mathscr{G}}_{s})]$ in terms of the time grid $\mathscr{G}$, where the expectation is taken over both system noise $W$ and the execution noises $(\xi_{i})_{i=0}^{\infty}$ (cf.  (3.1)).
4.1.1. Weak convergence for regular coefficients
To highlight the key ideas without introducing needless technicalities, we begin by providing the analysis for sufficiently smooth coefficients and test functions. In this setting, we obtain the maximum convergence rate, which is of the order $\mathcal{O}(|\mathscr{G}|)$ as $|\mathscr{G}|$ tends to zero within the regular time grids $\mathscr{P}_{[0,T]}$.
To specify the precise regularity conditions, we introduce the following notation: for each Euclidean space $E$ and $k\in{\mathbb{N}}$, $p>0$, let $C^{k}_{p}([0,T]\times{\mathbb{R}}^{d};E)$ be the space of functions $u:[0,T]\times{\mathbb{R}}^{d}\to E$ such that for all $r\in{\mathbb{N}}_{0}$ and multi-indices $s$ satisfying $2r+|s|\leq k$, the partial derivative $\partial_{t}^{r}\partial_{x}^{s}u$ exists and is continuous for all $(t,x)\in[0,T]\times\mathbb{R}^{d}$, and they all have polynomial growth in $x$:
$\|u\|_{C_{p}^{k}}:=\sum_{2r+|s|\leq k}\sup_{(t,x)\in[0,T]\times\mathbb{R}^{d}}\frac{\left|\partial_{t}^{r}\partial_{x}^{s}u(t,x)\right|}{1+|x|^{p}}<\infty.$ (4.1)
Let $C^{k,0}_{p}([0,T]\times{\mathbb{R}}^{d}\times A;E)$ be the space of functions $u:[0,T]\times{\mathbb{R}}^{d}\times{A}\to E$ such that for all $r\in{\mathbb{N}}_{0}$ and multi-indices $s$ satisfying $2r+|s|\leq k$, the partial derivative $\partial_{t}^{r}\partial_{x}^{s}u(t,x,a)$ exists, is continuous with respect to $t$ and $x$ for all $(t,x,a)\in[0,T]\times\mathbb{R}^{d}\times{A}$ and they all have polynomial growth in $x$ and $a$:
$\|u\|_{C_{p}^{k,0}}:=\sum_{2r+|s|\leq k}\sup_{(t,x,a)\in[0,T]\times\mathbb{R}^{d}\times A}\frac{\left|\partial_{t}^{r}\partial_{x}^{s}u(t,x,a)\right|}{1+|x|^{p}+d_{A}(a,a_{0})^{p}}<\infty,$ (4.2)
where $a_{0}\in A$ is the fixed element in (H.1). For simplicity, we omit the range space $E$ when no confusion arises. Moreover, we denote by $\mathcal{L}$ the generator of the aggregated dynamics (2.4) satisfying for all $\phi\in C^{2}([0,T]\times{\mathbb{R}}^{d})$,
$\mathcal{L}\phi(t,x)=\tilde{b}^{\bm{\pi}}(t,x)\nabla_{x}\phi(t,x)+\frac{1}{2}\operatorname*{Tr}\left((\tilde{\sigma}^{\bm{\pi}}(\tilde{\sigma}^{\bm{\pi}})^{\top})(t,x)\textrm{Hess}_{x}\phi(t,x)\right).$ (4.3)
The following assumption presents the regularity conditions to facilitate the analysis.
H. 3.
  1. (1)
    $b$ and $\sigma\sigma^{\top}$ are in the space ${C}^{2,0}_{p}([0,T]\times{\mathbb{R}}^{d}\times A)$;
  2. (2)
    $\tilde{b}^{\bm{\pi}},\tilde{\sigma}^{\bm{\pi}}$ are sufficiently regular, such that for any $f\in C_{p}^{4}(\mathbb{R}^{d})$ and any $t^{\prime}\in[0,T]$, the PDE
    $(\partial_{t}v+\mathcal{L}v)(t,x)=0,\quad(t,x)\in[0,t^{\prime})\times{\mathbb{R}}^{d};\quad v(t^{\prime},x)=f(x),\quad x\in{\mathbb{R}}^{d}$ (4.4)
    has a unique solution $v_{f}(\cdot,\cdot;t^{\prime})\in C_{p}^{4}([0,t^{\prime}]\times\mathbb{R}^{d})$. Moreover, $\|v_{f}(\cdot,\cdot;t^{\prime})\|_{C_{p}^{4}}\leq C\|f\|_{C_{p}^{4}}$, where $C$ is a constant depending only on $T,\tilde{b}^{\bm{\pi}},\tilde{\sigma}^{\bm{\pi}},p$.
Remark 4.1.
Condition (H.3) holds if the coefficients of the aggregated dynamics (2.4) are sufficiently regular. For instance, if $\tilde{b}^{\bm{\pi}}$ and $\tilde{\sigma}^{\bm{\pi}}$ are in the class $C_{p}^{4}$ and have bounded derivatives, then (H.3) can be verified using the probabilistic argument in [28].
Theorem 4.1.
Suppose (H.1), (H.2) and (H.3) hold with $p\geq 2$. Then for any $f\in C_{p}^{4}(\mathbb{R}^{d})$, there exists a constant $C_{p}\geq 0$ depending only on $T,\tilde{b}^{\bm{\pi}},\tilde{\sigma}^{\bm{\pi}},\bm{\pi},b,\sigma,p$ such that for all grids $\mathscr{G}$,
$\sup_{t\in[0,T]}\left|\mathbb{E}[f(X^{\mathscr{G}}_{t})]-\mathbb{E}^{W}[f(\widetilde{X}_{t})]\right|\leq C_{p}\|f\|_{C_{p}^{4}}|\mathscr{G}|.$ (4.5)
Proof.
See Section 7.1.1. ∎
For a given time grid $\mathscr{G}$, we execute the policy only at the nodes $t_{i}$, and use the same action over the time interval $[t_{i},t_{i+1})$ to obtain the sampled dynamics $X^{\mathscr{G}}$. Theorem 4.1 illustrates that with this policy execution method, for a given test function $f$, the error between $\mathbb{E}[f(X^{\mathscr{G}}_{t})]$ and the original aggregated dynamics $\mathbb{E}^{W}[f(\widetilde{X}_{t})]$ converges to 00 at the optimal first-order convergence $O(|\mathscr{G}|)$.
We illustrate several common cases in which the required assumptions in Theorem 4.1 hold.
Example 4.1 (General smooth coefficients).
Assume $\bm{\pi}(\mathrm{d}a|t,x)$ admits a density function with respect to a measure $\mu$ on $A$ (e.g., Lebesgue measure), i.e.,
$\bm{\pi}(\mathrm{d}a|t,x)=H(t,x,a)\mu(\mathrm{d}a).$ (4.6)
If it holds for any integer $r\geq 0$ and multi-index $s$ satisfying $2r+|s|\leq 4$ that
  1. ()
    For $\phi\in\{b,\sigma\}$, $(t,x,a)\mapsto\partial_{t}^{r}\partial_{x}^{s}\phi(t,x,a)$ is bounded and continuous.
  2. ()
    $(t,x)\mapsto\int_{A}|\partial_{t}^{r}\partial_{x}^{s}H(t,x,a)|\mu(\mathrm{d}a)$ is bounded.
  3. ()
    $\sigma\sigma^{\top}$ is uniformly elliptic.
Then $\tilde{b}^{\bm{\pi}}$ and $\tilde{\sigma}^{\bm{\pi}}$ are of class $C_{p}^{4}$ with bounded derivatives.
Example 4.2 (Gaussian policies).
Assume $A=\mathbb{R}^{m}$ and $\bm{\pi}(\cdot|t,x)$ is Gaussian with respect to the Lebesgue measure on $\mathbb{R}^{m}$, i.e.,
${\bm{\pi}(\mathrm{d}a|t,x)}=\varphi(a|\mu(t,x),\Sigma^{2}(t,x))\mathrm{d}a,$ (4.7)
where $\varphi(\cdot|\mu,\Sigma^{2})$ is the Gaussian density with mean $\mu$ and variance $\Sigma^{2}$. If the following conditions hold:
  • (1)
    For $\phi\in\{b,\sigma\}$, integer $r\geq 0$ and multi-index $s_{1},s_{2}$ satisfying that $0<2r+|s_{1}|+|s_{2}|\leq 4$, $(t,x,a)\mapsto\partial_{t}^{l}\partial_{x}^{s_{1}}\partial_{a}^{s_{2}}\phi(t,x,a)$ is bounded and continuous.
  • (2)
    $\mu,\Sigma\in C_{p}^{4}$ with bounded derivatives.
  • (3)
    $\sigma\sigma^{\top}$ is uniformly elliptic.
Then $\tilde{b}^{\bm{\pi}}$ and $\tilde{\sigma}^{\bm{\pi}}$ are of class $C_{p}^{4}$ with bounded derivatives.
Examples 4.1 and 4.2 assume that $\sigma\sigma^{\top}$ is uniformly elliptic, which implies the uniform ellipticity of $(t,x)\mapsto\int_{A}\sigma\sigma^{\top}(t,x,a)\bm{\pi}(\mathrm{d}a|t,x)$. This nondegeneracy condition ensures that after taking the square root, $(t,x)\mapsto\tilde{\sigma}^{\bm{\pi}}(t,x)$ has the same regularity as $(t,x)\mapsto\int_{A}\sigma\sigma^{\top}(t,x,a)\bm{\pi}(\mathrm{d}a|t,x)$. The existence of a sufficiently regular square root $\tilde{\sigma}^{\bm{\pi}}$ can be ensured by replacing the nondegeneracy condition with certain structural conditions on $\sigma$ and $\bm{\pi}$. Specifically, if we can express $\int_{A}\sigma\sigma^{\top}(t,x,a)\bm{\pi}(\mathrm{d}a|t,x)=\lambda(t,x)\lambda^{\top}(t,x)$ with a sufficiently smooth function $\lambda(t,x)$, then $\tilde{\sigma}^{\bm{\pi}}=\lambda$ will have the desired regularity, even if it is degenerate. We provide several examples to illustrate this.
Example 4.3 (Uncontrolled volatility).
Consider the case in which $\sigma$ is uncontrolled, i.e., $\sigma(t,x,a)=\sigma(t,x)$. In this case,
$\int_{A}\sigma\sigma^{\top}(t,x,a)\bm{\pi}(\mathrm{d}a|t,x)=\sigma\sigma^{\top}(t,x).$ (4.8)
Note that we only require $\tilde{\sigma}^{\bm{\pi}}(\tilde{\sigma}^{\bm{\pi}})^{\top}=\int_{A}\sigma\sigma^{\top}(t,x,a)\bm{\pi}(\mathrm{d}a|t,x)$, which means that we can choose $\tilde{\sigma}^{\bm{\pi}}$ to be $\sigma$ instead of $\sqrt{\sigma\sigma^{\top}}$. Therefore, the conditions on $\sigma$ in Examples 4.1 and 4.2 can be relaxed so that $\sigma$ is of class $C_{p}^{4}$ with bounded derivatives and it can be degenerate.
Example 4.4 (Gaussian policies with linear volatility).
Consider the case where the volatility linearly depends on the action, i.e., $\sigma(t,x,a)=a\beta^{\top}$ with $\beta\in R^{d}$ and (4.7) holds with $m=d$ and $\mu(t,x)=0$ and, then
$\int_{A}\sigma\sigma^{\top}(t,x,a)\bm{\pi}(\mathrm{d}a|t,x)=\int_{\mathbb{R}^{d}}\Sigma(t,x)u\beta^{\top}\beta u^{\top}\Sigma(t,x)^{\top}\varphi(u)\mathrm{d}u.$ (4.9)
We can choose $\tilde{\sigma}^{\bm{\pi}}(t,x)=\Sigma(t,x)\sqrt{\int_{\mathbb{R}^{d}}u\beta^{\top}\beta u^{\top}\varphi(u)\mathrm{d}u}$. Therefore, $\tilde{\sigma}^{\bm{\pi}}$ can degenerate; the assumption in this case is that $\Sigma\in C_{p}^{4}$ has bounded derivatives.
If $d=1$, then we can extend the second case to a general form. Assume that $\bm{\pi}(\cdot|t,x)$ is the location scaled distribution, i.e., there is a function $\Pi(t,x)$ such that $\bm{\pi}(\cdot|t,x)$ satisfies for any measurable function $f$, $\int_{A}f(a)\bm{\pi}(\mathrm{d}a|t,x)=\int_{B}f(\Pi(t,x)u)\mu(\mathrm{d}u)$, where $(B,\mathscr{F}_{B},\mu)$ is a measure space. Let $\sigma(t,x,a)=h(t,x)a$ for a bounded $h$. Moreover, assume that $h\geq 0$ and $\Pi\geq 0$. We have
$(\tilde{\sigma}^{\bm{\pi}}(t,x))^{2}=\int_{B}h^{2}(t,x)\Pi^{2}(t,x)u^{2}\mu(\mathrm{d}u).$ (4.10)
If we can choose $\tilde{\sigma}^{\bm{\pi}}(t,x)=h(t,x)\Pi(t,x)\sqrt{\int_{B}u^{2}\mu(\mathrm{d}u)}$, then $\tilde{\sigma}^{\bm{\pi}}$ can degenerate as long as $h\Pi\in C^{4}_{p}$ and has bounded derivatives.
4.1.2. Weak convergence with irregular coefficients
In the previous section, we established a first-order weak convergence of the sampled dynamics under the assumption that all coefficients are sufficiently regular so that the PDE associated with the aggregated dynamics has a four-times differentiable solution (see (H.3) and Remark 4.1). In this section, we relax the regularity conditions on the coefficients and quantify the weak convergence of the sampled dynamics based on their smoothness.
The following Hölder spaces will be used in the analysis.
Definition 4.1.
For $l\in(0,\infty)\setminus\mathbb{N}$, let $\mathcal{H}_{T}^{(l)}$ be the space of functions $u:[0,T]\times\mathbb{R}^{d}\times{A}\to\mathbb{R}$ such that for all integer $r\geq 0$ and multi-index $s$ satisfying $2r+|s|\leq l$, the partial derivatives $\partial_{t}^{r}\partial_{x}^{s}u$ are bounded and Hölder continuous with index $l-[l]$ in $x$ and with index $(l-2r-|s|)/2$ in $t$, i.e.,
$\|u\|_{T}^{(l)}:=$ $\sum_{2r+|s|\leq[l]}\sup_{t,x,a}|\partial_{t}^{r}\partial_{x}^{s}u(t,x,a)|$ (4.11)
$+\sum_{2r+|s|=[l]}\sup_{t,a,x\neq x^{\prime}}\frac{\left|\partial_{t}^{r}\partial_{x}^{s}u(t,x,a)-\partial_{t}^{r}\partial_{x}^{s}u(t,x^{\prime},a)\right|}{\left|x-x^{\prime}\right|^{l-[l]}}$
$+\sum_{0<l-2r-|s|<2}\sup_{t\neq t^{\prime},x,a}\frac{\left|\partial_{t}^{r}\partial_{x}^{s}u(t,x,a)-\partial_{t}^{r}\partial_{x}^{s}u(t^{\prime},x,a)\right|}{\left|t-t^{\prime}\right|^{(l-2r-|s|)/2}}<+\infty.$
Let $\mathcal{H}^{(l)}$ be the space of all functions $u\in\mathcal{H}_{T}^{(l)}$ that do not depend on $t$ and $a$, and let $\|\cdot\|^{(l)}$ be the corresponding norm.
The following regularity condition is imposed in this section.
H. 4.
Let $l\in(0,1)\cup(1,2)\cup(2,3)$, $b,\sigma\sigma^{\top}\in\mathcal{H}_{T}^{(l)}$, $\tilde{b}^{\bm{\pi}},\tilde{\sigma}^{\bm{\pi}}\in\mathcal{H}_{T}^{(l)}$, and $\tilde{\sigma}^{\bm{\pi}}(\tilde{\sigma}^{\bm{\pi}})^{\top}$ is uniformly elliptic.
Under (H.4), standard regularity theory for parabolic PDEs (see, e.g., [24, Chapter 4, Theorem 5.1]) shows that for any $f\in\mathcal{H}^{(l+2)}$ and $t^{\prime}\in[0,T]$, the PDE (4.4) has a unique solution $v_{f}(\cdot,\cdot;t^{\prime})\in\mathcal{H}_{t^{\prime}}^{(l+2)}$ and $\|v_{f}(\cdot,\cdot;t^{\prime})\|_{t^{\prime}}^{(l+2)}\leq C\|f\|^{(l+2)}$, where the constant $C$ depends only on $T,\tilde{b}^{\bm{\pi}}$ and $\tilde{\sigma}^{\bm{\pi}}$.
Leveraging this PDE regularity, the following theorem quantifies the weak convergence order of the sampled dynamics in terms of the smoothness coefficient $l$.
Theorem 4.2.
Suppose that (H.1), (H.2) and (H.4) hold. Then for $f\in\mathcal{H}^{(l+2)}$, there exists a constant $C$ depending only on $b,\sigma,\tilde{b}^{\bm{\pi}},\tilde{\sigma}^{\bm{\pi}},T$ such that for all grids $\mathscr{G}$,
$\sup_{t\in[0,T]}\left|\mathbb{E}\left[f(X^{\mathscr{G}}_{t})\right]-\mathbb{E}\left[f(\widetilde{X}_{t})\right]\right|\leq C\|f\|^{(l+2)}|\mathscr{G}|^{\mathcal{X}(l)},$ (4.12)
where the constant $\mathcal{X}(l)$ is defined by
$\mathcal{X}(l)\coloneqq\left\{\begin{aligned} &l/2,&l\in(0,1),\\ &1/(3-l),&l\in(1,2),\\ &1,&l\in(2,3).\end{aligned}\right.$ (4.13)
Proof.
See Section 7.1.2. ∎
Theorem 4.2 establishes the weak convergence order of the sampled dynamics while relaxing the regularity conditions on the coefficients in Theorem 4.1 and allowing for Hölder continuous coefficients. For clarity, we assume that the coefficients are bounded and that the volatility coefficients are uniformly elliptic, enabling us to leverage parabolic regularity theory to ensure the associated PDE admits a sufficiently regular classical solution. The result naturally extends to coefficients with polynomial growth and degenerate volatility, provided that the PDE (4.4) admits a classical solution with polynomial growth.
Here we emphasize that the analysis of policy execution error in Theorem 4.2 fundamentally differs from the convergence rate analysis of the classical piecewise constant control approximation in [16]. This distinction arises because the aggregated dynamics and the sampled dynamics have different coefficients and are thus associated with different PDEs—one governed by aggregated coefficients and the other by random coefficients influenced by execution noise. As a result, the viscosity solution arguments used in [16] are not directly applicable when the PDE (4.4) fails to admit a classical solution.
4.2. Weak error conditioned on execution noise
In this section, we establish high-probability bounds for the policy execution error with respect to the randomness used to sample the actions.
More precisely, let $\mathcal{F}^{\xi}:=\sigma\{\xi_{i}\mid{i\geq 0}\}$, and consider the error term $\mathbb{E}^{W}[f(X^{\mathscr{G}}_{t})]-\mathbb{E}^{W}[f(\widetilde{X}_{t})]$, where $f$ is a given test function, and $\mathbb{E}^{W}[\cdot]$ denotes the conditional expectation $\mathbb{E}[\cdot\mid\mathcal{F}^{\xi}]$. Note that the error $\mathbb{E}^{W}[f(X^{\mathscr{G}}_{t})]-\mathbb{E}^{W}[f(\widetilde{X}_{t})]$ is a random variable since it depends on the realization of the execution noises $\xi$. We have shown in Section 4.1 that the error converges to zero in expectation as $|\mathscr{G}|\to 0$. In the sequel, we provide a high-probability bound of this error.
H. 5.
For every $a\in A$, the functions $b(\cdot,\cdot,a)$ and $\sigma\sigma^{\top}(\cdot,\cdot,a)$ are of class $C_{p}^{2}([0,T]\times\mathbb{R}^{d})$ for some $p\geq 2$. Moreover, for $\varphi_{1}\in\{b,\sigma\}$, there is a constant $C\geq 0$ such that
$\left|\varphi_{1}(t,x,a)\right|\leq C\left(1+|x|\right),\quad\forall(t,x,a)\in[0,T]\times\mathbb{R}^{d}\times{A},$ (4.14)
and $\sup_{a\in A}\|b(\cdot,\cdot,a)\|_{C_{p}^{2}}<\infty$ and $\sup_{a\in A}\|\sigma\sigma^{\top}(\cdot,\cdot,a)\|_{C_{p}^{2}}<\infty$.
Remark 4.2.
Compared with (H.3), (H.5) imposes the additional condition on the uniform boundedness $\sup_{a\in A}\|b(\cdot,\cdot,a)\|_{C_{p}^{2}}<\infty$, and $\sup_{a\in A}\|\sigma\sigma^{\top}(\cdot,\cdot,a)\|_{C_{p}^{2}}<\infty$, which allows for controlling the concentration behavior of the sampled dynamics with respect to the execution noise $\xi$. It is easy to see that (H.4) implies (H.5).
The following theorem shows that the weak convergence rate in expectation, established in Section 4.1, also holds in a high probability sense with respect to the sampling noises. It extends a similar result established in [32] for LQ problems with Gaussian policies to general state dynamics and stochastic policies.
Theorem 4.3.
Suppose (H.1), (H.2), (H.3) and (H.5) hold. Then for the test function $f$ in Theorem 4.1, there is a constant $C_{p}$ depending only on $b,\sigma,\tilde{b}^{\bm{\pi}},\tilde{\sigma}^{\bm{\pi}},T,p$, such that for all grids $\mathscr{G}$ and all $\varrho\in(0,1)$,
$\sup_{t\in[0,T]}\mathbb{P}\left[\left|\mathbb{E}^{W}[f(X^{\mathscr{G}}_{t})]-\mathbb{E}^{W}[f(\widetilde{X}_{t})]\right|>C_{p}\|f\|_{C_{p}^{4}}\sqrt{|\mathscr{G}|}\big{(}1+\sqrt{\ln(2/\varrho)}\big{)}\right]\leq\varrho.$ (4.15)
Assume further that (H.1), (H.2) and (H.4) hold. Then for the test function $f$ in Theorem 4.2, there is a constant $C$ depending only on $b,\sigma,\tilde{b}^{\bm{\pi}},\tilde{\sigma}^{\bm{\pi}},T$ such that for all grids $\mathscr{G}$ and all $\varrho\in(0,1)$,
$\sup_{t\in[0,T]}\mathbb{P}\left[\left|\mathbb{E}^{W}[f(X^{\mathscr{G}}_{t})]-\mathbb{E}^{W}[f(\widetilde{X}_{t})]\right|>C\|f\|^{(l+2)}|\mathscr{G}|^{\mathcal{X}(l)\land\frac{1}{2}}\big{(}1+\sqrt{\ln(2/\varrho)}\big{)}\right]\leq\varrho,$ (4.16)
with $\mathcal{X}(l)$ defined by (4.13).
Proof.
See Section 7.1.3. ∎
The proof of Theorem 4.3, provided in Section 7.1.3, is based on establishing a pointwise upper bound for the weak error on each subinterval in terms of the grid size, and subsequently controlling the overall error using concentration inequalities for bounded martingale difference sequences.
4.3. Strong convergence
In this subsection, we quantify the pathwise difference between the sampled dynamics and the aggregated dynamics in terms of the sampling frequency for a fixed realization of the Brownian motion. To this end, we assume the volatility $\sigma$ is uncontrolled, i.e., $\sigma(t,x,a)=\sigma(t,x)$, and impose regularity conditions on the function $\phi$ used for sampling the actions (recall that $a_{t_{i}}=\phi(t_{i},X^{\mathscr{G}}_{t_{i}},\xi_{i})$ as in (3.2)).
H. 6.
  1. (1)
    There exists $C\geq 0$ such that for all $(t_{1},x_{1},a_{1}),(t_{2},x_{2},a_{2})\in[0,T]\times\mathbb{R}^{d}\times A$, $u\in E$,
    $\left|b(t_{1},x_{1},a_{1})-b(t_{2},x_{2},a_{2})\right|+d_{A}(\phi(t_{1},x_{1},u),\phi(t_{2},x_{2},u))\leq C\left(\left|t_{1}-t_{2}\right|^{\frac{1}{2}}+\left|x_{1}-x_{2}\right|+d_{A}(a_{1},a_{2})\right),$ (4.17)
    and for all $x,y\in\mathbb{R}^{d}$ and $t\in[0,T]$,
    $\left|\sigma(t,x)-\sigma(t,y)\right|\leq C\left|x-y\right|.$ (4.18)
  2. (2)
    There exists $(t_{0},x_{0},a_{0})\in[0,T]\times\mathbb{R}^{d}\times A$, such that $d_{A}(\phi(t_{0},x_{0},\xi),a_{0})$ is sub-exponential  [^3] We say a random variable $X$ is sub-exponential if there exists a constant $C>0$ such that $\mathbb{E}[|X|>t]\leq 2e^{-Ct}$ for every $t\geq 0$..
Example 4.5.
If $\bm{\pi}(\cdot|t,x)$ satisfies (4.7) with $\Sigma(t,x)=\Sigma$, then we can write $a=\phi(t,x,\xi)=\mu(t,x)+\Sigma\xi$. In this case, the assumptions for $\phi$ in (H.6) hold. This corresponds to the commonly used Gaussian policies considered in [32] and many others.
Theorem 4.4.
Suppose (H.2) and (H.6) hold. For all $p\geq 2$ and all grids $\mathscr{G}=\{0=t_{0}<\ldots<t_{n}=T\}$ with $|\mathscr{G}|^{1/2}\log(1+n)\leq C_{G}<\infty$,
$\mathbb{E}\left[\sup_{0\leq t\leq T}\left|X^{\mathscr{G}}_{t}-\widetilde{X}_{t}\right|^{p}\right]\leq C|\mathscr{G}|^{\frac{p}{2}},$ (4.19)
where the constant $C$ depends only on $b,\sigma,\phi,T,p$ and $C_{G}$.
Consequently, if $\mathscr{G}=\{t_{i}\}_{i=0}^{n}$ with $t_{i}=i\frac{T}{n}$, then for every $\varepsilon>0$ and almost every $\omega\in\Omega$, there is an integer $N_{\varepsilon}(\omega)$ such that for every $n\geq N_{\varepsilon}(\omega)$,
$\sup_{0\leq t\leq T}\left|X^{\mathscr{G}}_{t}(\omega)-\widetilde{X}_{t}(\omega)\right|\leq n^{-\frac{1}{2}+\varepsilon}.$ (4.20)
The proof of Theorem 4.4 is given in Section 7.1.4. It involves carefully estimating the Orlicz norm of the pathwise difference between the aggregated and sampled dynamics on each subinterval and subsequently controlling the overall error using concentration inequalities for sub-exponential martingale difference sequences.
Theorem 4.4 indicates that the convergence of the sampled dynamics to the aggregated dynamics can be interpreted as a law of large numbers in time. Specifically, when the volatility is uncontrolled, the $1/2$-order convergence is achieved by increasing the sampling frequency for any fixed realization of the system noise.
Unfortunately, such a strong convergence fails when the volatility is controlled, as shown in the following example.
Example 4.6.
Let $d=1$, $A\subset{\mathbb{R}}$, $b(t,x,a)=0$ and $\sigma(t,x,a)=a$. Assume $\bm{\pi}(\mathrm{d}a|t,x)=\mu(\mathrm{d}a)$ for some $\mu\in\mathcal{P}(A)$ and $Q:=\int_{A}a^{2}\mu(\mathrm{d}a)<+\infty$. Assume that $\{\xi_{i}\}_{i=0}^{\infty}$ are i.i.d. random variables with distribution $\mu$. Let $a_{t_{i}}=\xi_{i}$ for all $i$, then we have
$X^{\mathscr{G}}_{t}-\widetilde{X}_{t}=\int_{0}^{t}\left(a_{\delta(s)}-Q^{1/2}\right)\mathrm{d}W_{s}.$ (4.21)
Note that $W$ is a Brownian motion under the filtration $\{\mathcal{F}_{t}\}_{t\geq 0}$ with $\mathcal{F}_{t}:=\sigma\{a_{\delta(s)}:s\leq t\}\lor\mathcal{F}^{W}_{t}$. It follows from Itô’s isometry that
$\mathbb{E}\left[|X^{\mathscr{G}}_{t}-\widetilde{X}_{t}|^{2}\right]=\mathbb{E}\left[\int_{0}^{t}(a_{\delta(s)}-Q^{1/2})^{2}\mathrm{d}s\right]=t\mathbb{E}\left[(\xi_{1}-Q^{1/2})^{2}\right],$ (4.22)
which does not converge to zero as $|\mathscr{G}|\rightarrow 0$, except that $\xi_{1}$ is a constant (or equivalently $\mu$ is a Dirac distribution).
5. Applications
This section builds on the convergence results from Section 4 to quantify the accuracy of various estimators used in continuous-time RL algorithms. The proofs of all statements are given in Section 7.2.
5.1. Accuracy of the Monte-Carlo evaluation
Recall that in RL, the agent has the objective function (2.2), and she needs to estimate its value under a given policy $\bm{\pi}$ with data that is generated by executing this policy. This problem is known as the (on-policy) policy evaluation.
The agent can only discretely sample the actions at the pre-specified time grid and observe the state variable at the same time grid $\mathscr{G}=\{0=t_{0}<\ldots<t_{n}=T\}$. Therefore, the agent can sample the process $X^{\mathscr{G}}_{s}$. For simplicity, we assume that the agent can collect and observe the accumulated running reward $R_{t_{i}}^{\mathscr{G}}=\int_{t_{i}}^{t_{i+1}}r(s,X^{\mathscr{G}}_{s},a_{t_{i}})\mathrm{d}s$ [^4] If only the instantaneous reward $r(s,X^{\mathscr{G}}_{t_{i-1}},a_{t_{i}})$ can be observed at the grid point, then the accumulated reward need to be further approximated as $R_{t_{i}}^{\mathscr{G}}\approx r(s,X^{\mathscr{G}}_{t_{i-1}},a_{t_{i}})(t_{i}-t_{i-1})$ as in the usual Euler scheme. Under suitable assumptions on the regularity conditions on the function $r$, the order of our analysis does not change. (cf. [18]) Hence, the expected reward collected by the agent is
$J^{\mathscr{G}}\coloneqq\mathbb{E}\left[\int_{0}^{T}r(s,X^{\mathscr{G}}_{s},a_{\delta(s)})\mathrm{d}s+h(X^{\mathscr{G}}_{T})\right]=\mathbb{E}\left[\sum_{i=0}^{n-1}R_{t_{i}}^{\mathscr{G}}+h(X^{\mathscr{G}}_{T})\right].$ (5.1)
In, the expected reward with a continuous-time execution of the policy is given by (2.6):
$\tilde{J}\coloneqq\mathbb{E}\left[\int_{0}^{T}\tilde{r}(s,\widetilde{X}_{s})\mathrm{d}s+h(\widetilde{X}_{T})\right]=\mathbb{E}\left[\int_{0}^{T}\int_{A}r(s,\widetilde{X}_{s},a)\bm{\pi}(\mathrm{d}a|s,\widetilde{X}_{s})\mathrm{d}s+h(\widetilde{X}_{T})\right].$ (5.2)
The following corollary characterizes the difference between (5.2) and (5.1), that is, the bias of using $\int_{0}^{T}r(s,X^{\mathscr{G}}_{s},a_{\delta(s)})\mathrm{d}s+h(X^{\mathscr{G}}_{T})$ as an estimate for $\tilde{J}$.
Proposition 5.1.
Assume (H.1) and (H.2) hold.
  1. (1)
    If $r\in C^{2,0}_{p},\tilde{r}\in C^{2}_{p}$ and (H.3) holds with the terminal function $f=h,\tilde{r}(t,\cdot),0\leq t\leq T$, then there is a constant $C_{p}$ depending only on $b,\sigma,\tilde{b}^{\bm{\pi}},\tilde{\sigma}^{\bm{\pi}},\bm{\pi},T,p$ such that
    $\left|J^{\mathscr{G}}-\tilde{J}\right|\leq C_{p}\left(\|h\|_{C_{p}^{4}}+\|r\|_{C_{p}^{2,0}}+\sup_{t\in[0,T]}\|\tilde{r}(t,\cdot)\|_{C_{p}^{4}}+\|\tilde{r}\|_{C_{p}^{2}}\right)|\mathscr{G}|.$ (5.3)
  2. (2)
[…]
    $\left|J^{\mathscr{G}}-\tilde{J}\right|\leq C\left(\|h\|^{(l+2)}+\|r\|_{T}^{(l)}+\sup_{t\in[0,T]}\|\tilde{r}(t,\cdot)\|^{(l+2)}+\|\tilde{r}\|_{T}^{(l)}\right)|\mathscr{G}|^{\mathcal{X}(l)},$ (5.4)
    with $\mathcal{X}(l)$ defined by (4.13).
Proposition 5.1 characterizes the (unconditional) bias of the estimated value of a given policy by discrete sampling and piecewise constant execution. Given this result, one may understand the required simulation budget in order to obtain an estimate with desired accuracy. Next, we use the notion of the conditional weak error to highlight that the required simulation budget can be refined to those conditioned on the policy sampler. More precisely, consider the value conditional on $\xi$,
$J^{\mathscr{G}}(\xi)\coloneqq\mathbb{E}^{W}\left[\int_{0}^{T}r(s,X^{\mathscr{G}}_{s},a_{\delta(s)})\mathrm{d}s+h(X^{\mathscr{G}}_{T})\right].$ (5.5)
Combining the proof of Proposition 5.1 and Theorem 4.3, we obtain the following result for the large deviation bound for the value function.
Proposition 5.2.
Assume (H.1) and (H.2) hold. If (H.5) and the conditions in Proposition 5.1(1) hold, then there is a constant $C_{p}$ depending only on $b,\sigma,\tilde{b}^{\bm{\pi}},\tilde{\sigma}^{\bm{\pi}},T,p$ such that for every $\varrho\in(0,1)$,
$\mathbb{P}\left[|J^{\mathscr{G}}(\xi)-\tilde{J}|>C_{p}\left(\|h\|_{C_{p}^{4}}+\|r\|_{C_{p}^{2,0}}+\sup_{t\in[0,T]}\|\tilde{r}(t,\cdot)\|_{C_{p}^{4}}+\|\tilde{r}\|_{C_{p}^{2}}\right)\sqrt{|\mathscr{G}|}(1+\sqrt{\ln(2/\varrho)})\right]\leq\varrho.$ (5.6)
If conditions in Proposition 5.1(2) holds, then there is a constant $C$ depending only on $b,\sigma,\tilde{b}^{\bm{\pi}},\tilde{\sigma}^{\bm{\pi}},T$ such that for every $\varrho\in(0,1)$,
$\mathbb{P}\left[|J^{\mathscr{G}}(\xi)-\tilde{J}|>C\left(\|h\|^{(l+2)}+\|r\|_{T}^{(l)}+\sup_{t\in[0,T]}\|\tilde{r}(t,\cdot)\|^{(l+2)}+\|\tilde{r}\|_{T}^{(l)}\right)|\mathscr{G}|^{\mathcal{X}(l)\land\frac{1}{2}}(1+\sqrt{\ln(2/\varrho)})\right]\leq\varrho$ (5.7)
with $\mathcal{X}(l)$ defined by (4.13).
5.2. Accuracy of Temporal-Difference learning
The temporal difference (TD) plays a crucial role in many RL algorithms as the policy evaluation methods. In contrast to the Monte-Carlo evaluation, TD learning often aims to find an approximation of the value function associated with a given policy $\bm{\pi}$, defined as
$\tilde{J}(t,x;\bm{\pi})=\mathbb{E}\left[\int_{t}^{T}\tilde{r}(s,\widetilde{X}_{s})\mathrm{d}s+h(\widetilde{X}_{T})\Big{|}\widetilde{X}_{t}=x\right],$ (5.8)
within a suitable parametric family of functions, and it makes use of the increment of the value function process.
In particular, suppose $V^{\theta}$ is a family of parameterized functions of $(t,x)$ with $\theta$ a finite-dimensional parameter. TD learning concerns the increment in the form “$\mathrm{d}V^{\theta}(t,\widetilde{X}_{t})$” and the instantaneous reward “$r(t,\widetilde{X}_{t})\mathrm{d}t$”. In the continuous-time framework, these two increments have different natures. “$\mathrm{d}V^{\theta}(t,\widetilde{X}_{t})$” is merely an informal notation that shall be combined with a suitable integrand as in the definition of Itô’s integral, and the latter “$r(t,\widetilde{X}_{t})\mathrm{d}t$” is the usual Lebesgue integral with respect to time. In theory, [19] identify the continuous-time counterpart of TD learning as solving the martingale orthogonality condition. More precisely, if $\theta^{*}$ is the root to
$\mathbb{E}^{W}\left[\int_{0}^{T}S(t,\widetilde{X}_{t})\left[\mathrm{d}V^{\theta}(t,\widetilde{X}_{t})+\tilde{r}(t,\widetilde{X}_{t})\mathrm{d}t\right]\right]=0,$ (5.9)
for all admissible test functions $S(t,x)$, then the $V^{\theta^{*}}(t,x)=\tilde{J}(t,x;\bm{\pi})$ is the desired value function associated with a given policy $\bm{\pi}$. Therefore, various RL algorithms are based on using stochastic approximation (cf. [7]) to find the root $\theta^{*}$ to the martingale orthogonality condition (5.9). The most popular choice is the so-called “TD(0)” algorithm, which corresponds to taking $S(t,x)=\frac{\partial V^{\theta}}{\partial\theta}(t,x)$.
For a given test function $S$ and parameter $\theta$, applying the stochastic approximation algorithm means approximating this expectation with random samples, i.e., an estimator of this expectation. The typical convergence conditions of the stochastic approximation algorithm are threefold [^5] Besides these conditions on the accuracy, the convergence also depends on the landscape of (5.9) as a function of $\theta$. However, this is jointly determined by the parametrization $\theta\mapsto V^{\theta}$ and the properties of the aggregated dynamics, which is not affected by time discretization.: first, the estimator has diminishing bias; second, the estimator has non-exploding variances; third, the learning rates should be suitably chosen according to the bias and variance of the estimator. Therefore, it is crucial to understand the magnitude of bias and variance and how they depend on the grid size.
In continuous time, however, due to the discretely sampled actions and the nature of the stochastic integral inside (5.9), the agent only has access to the sampled dynamics and uses the discrete sum (we drop the superscript $\theta$ in the following to illustrate the results for a generic function $V$)
$S\circ V:=\sum_{i=0}^{n-1}S(t_{i},X^{\mathscr{G}}_{t_{i}})\left(V(t_{i+1},X^{\mathscr{G}}_{t_{i+1}})-V(t_{i},X^{\mathscr{G}}_{t_{i}})+R_{t_{i}}^{\mathscr{G}}\right),$
with the observed accumulated reward $R_{t_{i}}^{\mathscr{G}}=\int_{t_{i}}^{t_{i+1}}r(s,X^{\mathscr{G}}_{s},a_{t_{i}})\mathrm{d}s$ as an approximation for the expectation in (5.9). We characterize the bias and variance of such an approximation in Proposition 5.3.
Proposition 5.3.
Suppose (H.1) and (H.2) hold. Assume that $V\in C_{p}^{2}$. Denote
$g(t,x,a)=\partial_{t}V(t,x)+b^{\top}(t,x,a)\partial_{x}V(t,x)+\frac{1}{2}\operatorname*{Tr}\left(\sigma\sigma^{\top}(t,x,a)\textrm{Hess}_{x}V(t,x)\right)+r(t,x,a)=:\mathcal{L}^{a}V(t,x).$ (5.10)
For the variance of $S\circ V$, if $S\in C_{p}^{0}$ and $r\in C^{0,0}_{p}$, then
$\mathrm{Var}\left(S\circ V\right)\leq C_{p}\|S\|_{C_{p}^{0}}^{2}(\|V\|_{C_{p}^{2}}+\|r\|_{C_{p}^{0,0}})^{2},$ (5.11)
where $C_{p}$ is a constant depending only on $b,\sigma,\tilde{b}^{\bm{\pi}},\tilde{\sigma}^{\bm{\pi}},\bm{\pi},T,p$.
For the bias of $S\circ V$, we have the following holds.
  1. (1)
    If $g\in C_{p}^{0,0},S\in C^{2}_{p}$ and $gS$ satisfies the condition (1) in Proposition 5.1 as the running term, then there is a constant $C_{p}$ depending only on $b,\sigma,\tilde{b}^{\bm{\pi}},\tilde{\sigma}^{\bm{\pi}},\bm{\pi},T,p$ such that
    $\left|\mathbb{E}[S\circ V]-\mathbb{E}^{W}\left[\int_{0}^{T}S(t,\widetilde{X}_{t})\left[\mathrm{d}V(t,\widetilde{X}_{t})+\tilde{r}(t,\widetilde{X}_{t})\mathrm{d}t\right]\right]\right|$ (5.12)
    $\leq C_{p}\left(\|S\|_{C_{p}^{2}}\|g\|_{C_{p}^{0,0}}+\|Sg\|_{C_{p}^{2,0}}+\|\widetilde{Sg}\|_{C_{p}^{2}}+\sup_{t\in[0,T]}\|\widetilde{Sg}(t,\cdot)\|_{C_{p}^{4}}\right)|\mathscr{G}|,$
    where $\widetilde{Sg}(t,x):=\int_{A}Sg(t,x,a)\bm{\pi}(\mathrm{d}a|t,x)$.
  2. (2)
    If $g\in\mathcal{H}_{T}^{(0)},S\in\mathcal{H}_{T}^{(l)}$ and $gS$ satisfies the condition (2) in Proposition 5.1 as the running term, then there is a constant $C$ depending only on $b,\sigma,\tilde{b}^{\bm{\pi}},\tilde{\sigma}^{\bm{\pi}},T$ such that
    $\left|\mathbb{E}[S\circ V]-\mathbb{E}^{W}\left[\int_{0}^{T}S(t,\widetilde{X}_{t})\left[\mathrm{d}V(t,\widetilde{X}_{t})+\tilde{r}(t,\widetilde{X}_{t})\mathrm{d}t\right]\right]\right|$ (5.13)
    $\leq C\left(\|S\|_{T}^{(l)}\|g\|_{T}^{(0)}+\|Sg\|^{(l)}_{T}+\|\widetilde{Sg}\|^{(l)}_{T}+\sup_{t\in[0,T]}\|\widetilde{Sg}(t,\cdot)\|^{(l+2)}\right)|\mathscr{G}|^{\mathcal{X}(l)}.$
Proposition 5.3 implies that we can obtain a more accurate stochastic approximation with a refined grid. In addition, the variance of such approximate is uniformly bounded in the grid size. It means that with a suitable schedule to refine the grid, we may expect the typical convergence conditions in the stochastic approximation algorithm will be satisfied.
5.3. Accuracy of policy gradient
The policy gradient is one of the most commonly used methods in policy optimization in RL. The well-celebrated (model-free) policy gradient representation (e.g., [31]) gives an unbiased estimator for the gradient of the value function with respect to certain parameters in the policy. For example, the stochastic policy $\bm{\pi}^{\psi}$ is parameterized by $\psi$, a finite-dimensional parameter, and it admits a density function with respect to the Lebesgue measure. We denote the density function by $H^{\psi}(t,x,a)=\frac{\bm{\pi}^{\psi}(\mathrm{d}a|t,x)}{\mu(\mathrm{d}a)}$. Then we know the value function under the policy $\bm{\pi}^{\psi}$ (defined in (5.8)) will depend on $\psi$ implicitly, denoted by $\tilde{J}(t,x;\bm{\pi}^{\psi})$. The policy gradient concerns about the derivative of $\tilde{J}(t,x;\bm{\pi}^{\psi})$ with respect to $\psi$, i.e., $\frac{\partial\tilde{J}(t,x;\bm{\pi}^{\psi})}{\partial\psi}$. For simplicity, we focus on the initial time-state gradient and suppress the time-state pair’s dependence.
A heuristic derivation in [19] shows that the gradient $\frac{\partial\tilde{J}(\bm{\pi}^{\psi})}{\partial\psi}$ can be represented by
$\mathbb{E}\left[\int_{0}^{T}\frac{\partial\log H^{\psi}}{\partial\psi}(a_{t}|t,X_{t})\left[\mathrm{d}\tilde{J}(t,X_{t};\bm{\pi}^{\psi})+r(t,X_{t},a_{t})\mathrm{d}t\right]\right]$
$=\mathbb{E}^{W}\left[\int_{0}^{T}\int_{A}\frac{\partial\log H^{\psi}}{\partial\psi}(a|t,\widetilde{X}_{t})\mathcal{L}^{a}\tilde{J}(t,\widetilde{X}_{t};\bm{\pi}^{\psi})\bm{\pi}(\mathrm{d}a|t,\widetilde{X}_{t})\mathrm{d}t\right].$ (5.14)
The expectation (5.14) can be approximated by the forward-Euler-type of finite sums using sampled dynamics, i.e.,
$\sum_{i=0}^{n-1}\frac{\partial\log H^{\psi}}{\partial\psi}(a_{t_{i}}|t_{i},X^{\mathscr{G}}_{t_{i}})\left[\tilde{J}(t_{i+1},X^{\mathscr{G}}_{t_{i+1}};\bm{\pi}^{\psi})-\tilde{J}(t_{i},X^{\mathscr{G}}_{t_{i}};\bm{\pi}^{\psi})+R_{t_{i}}^{\mathscr{G}}\right].$
Naturally, the convergence of gradient-based algorithms depends on whether we have an accurate gradient estimate. To analyze the accuracy of the policy gradient estimator, we consider the following general expression:
$S\circ V:=\sum_{i=0}^{n-1}S(t_{i},X^{\mathscr{G}}_{t_{i}},a_{t_{i}})\left(V(t_{i+1},X^{\mathscr{G}}_{t_{i+1}})-V(t_{i},X^{\mathscr{G}}_{t_{i}})+R_{t_{i}}^{\mathscr{G}}\right),$ (5.15)
where $R_{t_{i}}^{\mathscr{G}}=\int_{t_{i}}^{t_{i+1}}r(s,X^{\mathscr{G}}_{s},a_{t_{i}})\mathrm{d}s$ is the observed accumulated reward over $[t_{i},t_{i+1}]$.
The form (5.15) also appears in the continuous-time q-learning algorithm proposed in [20], where $S$ is the derivative of the q-function that needs to be learned. The difference between (5.15) and (5.9) is that now the function $S$ also depends on the sampled action $a_{t_{i}}$. Thus, (5.15) is not a straightforward substitution and discretization of any integration along the aggregated dynamics and can only be well-defined using the sampled dynamics. Note that given $X^{\mathscr{G}}_{t_{i}}$, the value of $X^{\mathscr{G}}_{t_{i+1}}$ not only depends on the environmental noise (the underlying Brownian motions) but also depends on $a_{t_{i}}$. The subtle difference in this measurability condition will affect the limiting value of (5.15).
We characterize the limit of a general estimator in the form (5.15) and quantify the magnitude of its bias and variance in the next proposition. The proof is similar to that of Proposition 5.3, so we omit it.
Proposition 5.4.
Suppose (H.1) and (H.2) hold. Assume $V\in C_{p}^{2}$. Define $g$ by (5.10). For the variance of $S\circ V$, if $S\in C_{p}^{0,0}$ and $r\in C_{p}^{0,0}$, then
$\mathrm{Var}\left(S\circ V\right)\leq C_{p}\|S\|_{C_{p}^{0,0}}^{2}(\|V\|_{C_{p}^{2}}+\|r\|_{C_{p}^{0,0}})^{2},$ (5.16)
where $C_{p}$ is a constant depending only on $b,\sigma,\tilde{b}^{\bm{\pi}},\tilde{\sigma}^{\bm{\pi}},\bm{\pi},T,p$.
For the bias of $S\circ V$, we have the following holds.
  1. (1)
    If $g\in C_{p}^{0,0},S\in C^{2,0}_{p}$ and $gS$ satisfies the condition (1) in Proposition 5.1 as the running term, then there is a constant $C_{p}$ depending only on $b,\sigma,\tilde{b}^{\bm{\pi}},\tilde{\sigma}^{\bm{\pi}},\bm{\pi},T,p$ such that
    $\left|\mathbb{E}[S\circ V]-\mathbb{E}^{W}\left[\int_{0}^{T}\widetilde{Sg}(t,\widetilde{X}_{t})\mathrm{d}t\right]\right|$ (5.17)
    $\leq C_{p}\left(\|S\|_{C_{p}^{2,0}}\|g\|_{C_{p}^{0,0}}+\|Sg\|_{C_{p}^{2,0}}+\|\widetilde{Sg}\|_{C_{p}^{2}}+\sup_{t\in[0,T]}\|\widetilde{Sg}(t,\cdot)\|_{C_{p}^{4}}\right)|\mathscr{G}|,$
    where $\widetilde{Sg}(t,x):=\int_{A}S(t,x,a)g(t,x,a)\bm{\pi}(\mathrm{d}a|t,x)$.
    In particular, if the conditions in [19] are satisfied so that the policy gradient has the representation (5.14), then by setting $V(\cdot,\cdot)=\tilde{J}(\cdot,\cdot;\bm{\pi}^{\psi})$, and $S(t,x,a)=\frac{\partial\log H^{\psi}}{\partial\psi}(a|t,x)$, we have
    $\left|\mathbb{E}\left[S\circ V\right]-\frac{\partial\tilde{J}(\bm{\pi}^{\psi})}{\partial\psi}\right|\leq C_{p}\left(\|S\|_{C_{p}^{2,0}}\|g\|_{C_{p}^{0,0}}+\|Sg\|_{C_{p}^{2,0}}+\|\widetilde{Sg}\|_{C_{p}^{2}}+\sup_{t\in[0,T]}\|\widetilde{Sg}(t,\cdot)\|_{C_{p}^{4}}\right)|\mathscr{G}|.$
  2. (2)
    If $g\in\mathcal{H}_{T}^{(0)},S\in\mathcal{H}_{T}^{(l)}$ and $gS$ satisfies the condition (2) in Proposition 5.1 as the running term, then there is a constant $C$ depending only on $b,\sigma,\tilde{b}^{\bm{\pi}},\tilde{\sigma}^{\bm{\pi}},T$ such that
    $\left|\mathbb{E}[S\circ V]-\mathbb{E}^{W}\left[\int_{0}^{T}\widetilde{Sg}(t,\widetilde{X}_{t})\mathrm{d}t\right]\right|$ (5.18)
    $\leq C\left(\|S\|_{T}^{(l)}\|g\|_{T}^{(0)}+\|Sg\|^{(l)}_{T}+\|\widetilde{Sg}\|^{(l)}_{T}+\sup_{t\in[0,T]}\|\widetilde{Sg}(t,\cdot)\|^{(l+2)}\right)|\mathscr{G}|^{\mathcal{X}(l)}.$
    In particular, if the conditions in [19] are satisfied so that the policy gradient has the representation (5.14), then by setting $V(\cdot,\cdot)=\tilde{J}(\cdot,\cdot;\bm{\pi}^{\psi})$, and $S(t,x,a)=\frac{\partial\log H^{\psi}}{\partial\psi}(a|t,x)$, we have
    $\left|\mathbb{E}\left[S\circ V\right]-\frac{\partial\tilde{J}(\bm{\pi}^{\psi})}{\partial\psi}\right|\leq C\left(\|S\|_{T}^{(l)}\|g\|_{T}^{(0)}+\|Sg\|^{(l)}_{T}+\|\widetilde{Sg}\|^{(l)}_{T}+\sup_{t\in[0,T]}\|\widetilde{Sg}(t,\cdot)\|^{(l+2)}\right)|\mathscr{G}|^{\mathcal{X}(l)}.$
Proposition 5.4 implies that with a refined grid, we can obtain a more accurate estimate for the policy gradient. In addition, the variance of such an estimator is uniformly bounded in the grid size. These properties are crucial for adjusting the sampling frequency to ensure the convergence of the overall policy gradient algorithm. [^6] The typical conditions of convergence of the stochastic gradient descent algorithm also require $\frac{\partial\tilde{J}(\bm{\pi}^{\psi})}{\partial\psi}$ to satisfy certain conditions as a function of $\psi$, but it is not affected by time discretization. Note that in a typical actor-critic algorithm, one does not have an exact estimate of $\tilde{J}(\cdot,\cdot;\bm{\pi}^{\psi})$, and can only obtain its approximation $V^{\theta}\approx\tilde{J}(\cdot,\cdot;\bm{\pi}^{\psi})$ via policy evaluation. Therefore, the convergence of a general actor-critic algorithm is more complex and open for future research.
5.4. Accuracy of risk-sensitive q-learning
So far, all the discussions have been restricted to the additive functional, which is often regarded as a risk-neutral problem. In continuous-time risk-sensitive problems (e.g., [6]), one encounters the objective function in the exponential form:
$\frac{1}{\epsilon}\log\mathbb{E}\left[\exp\left\{\epsilon\left[\int_{0}^{T}r(s,X_{s},a_{s})\mathrm{d}s+h(X_{T})\right]\right\}\right].$
[17] establishes the risk-sensitive q-learning theory and reveals that it only differs from the non-risk-sensitive counterpart by a penalty on the quadratic variation of the value function. Hence, the RL algorithm for the non-risk-sensitive problems can be easily modified to account for risk sensitivity by adding an integral with respect to a suitable quadratic variation process, which can be approximated by the sampled dynamics as
$S\Box V:=\sum_{i=0}^{n-1}S(t_{i},X^{\mathscr{G}}_{t_{i}},a_{t_{i}})\left(V(t_{i+1},X^{\mathscr{G}}_{t_{i+1}})-V(t_{i},X^{\mathscr{G}}_{t_{i}})+R_{t_{i}}^{\mathscr{G}}\right)^{2}.$
Proposition 5.5.
Suppose (H.1) and (H.2) hold. Assume $V\in C_{p}^{2}$. Define $g_{1}$ by (5.10) and $g_{2}(t,x,a)=\left|\sigma^{\top}\partial_{x}V(t,x,a)\right|^{2}$. For the variance of $S\circ V$, if $S,g_{1}\in C_{p}^{0,0}$, then
$\mathrm{Var}\left(S\Box V\right)\leq C_{p}\|S\|_{C_{p}^{0,0}}^{2}\left(\|g_{1}\|_{C_{p}^{0,0}}^{2}+\|V\|_{C_{p}^{2}}^{2}\right)^{2},$ (5.19)
where $C_{p}$ is a constant depending only on $b,\sigma,\tilde{b}^{\bm{\pi}},\tilde{\sigma}^{\bm{\pi}},\bm{\pi},T,p$.
For the bias of $S\Box V$, we have the following holds.
  1. (1)
    If $S,g_{1},\partial_{x}V,\sigma\in C^{2,0}_{p},g_{2}\in C_{p}^{0,0}$ and $Sg_{2}$ satisfies the condition (1) on Proposition 5.1 as the running term, then there is a constant $C_{p}$ depending only on $b,\sigma,\tilde{b}^{\bm{\pi}},\tilde{\sigma}^{\bm{\pi}},\bm{\pi},T,p$ such that
    $\left|\mathbb{E}[S\Box V]-\mathbb{E}\left[\int_{0}^{T}\widetilde{Sg_{2}}(t,\widetilde{X}_{t})\mathrm{d}t\right]\right|$ (5.20)
    $\leq C_{p}\left(\|S\|_{C_{p}^{2,0}}(\|g_{1}\|_{C_{p}^{2,0}}^{2}+\|g_{2}\|_{C_{p}^{0,0}})+\|g_{1}\|_{C_{p}^{2,0}}\|V\|_{C_{p}^{2}}+\|Sg_{2}\|_{C_{p}^{2,0}}+\|\widetilde{Sg_{2}}\|_{C_{p}^{2}}+\sup_{t\in[0,T]}\|\widetilde{Sg_{2}}(t,\cdot)\|_{C_{p}^{4}}\right)|\mathscr{G}|.$
  2. (2)
    If $S,g_{1},\partial_{x}V,\sigma\in\mathcal{H}^{(l)}_{T},g_{2}\in\mathcal{H}_{T}^{(0)}$ and $Sg_{2}$ satisfies the condition (2) on Proposition 5.1 as the running term, then there is a constant $C$ depending only on $b,\sigma,\tilde{b}^{\bm{\pi}},\tilde{\sigma}^{\bm{\pi}},T$ such that
    $\left|\mathbb{E}[S\Box V]-\mathbb{E}\left[\int_{0}^{T}\widetilde{Sg_{2}}(t,\widetilde{X}_{t})\mathrm{d}t\right]\right|$ (5.21)
    $\leq C\left(\|S\|_{T}^{(l)}((\|g_{1}\|_{T}^{(l)})^{2}+\|g_{2}\|_{T}^{(0)})+\|g_{1}\|_{T}^{(l)}\|V\|_{T}^{(l)}+\|Sg_{2}\|_{T}^{(l)}+\|\widetilde{Sg_{2}}\|_{T}^{(l)}+\sup_{t\in[0,T]}\|\widetilde{Sg_{2}}(t,\cdot)\|^{(l+2)}\right)|\mathscr{G}|^{\mathcal{X}(l)}.$
In particular, if $S$ does not depend on $a$, then
$\mathbb{E}\left[\int_{0}^{T}\int_{A}\widetilde{Sg_{2}}(t,\widetilde{X}_{t})\mathrm{d}t\right]=\mathbb{E}\left[\int_{0}^{T}S(t,\widetilde{X}_{t})\mathrm{d}\left<V(\cdot,\widetilde{X}_{\cdot})\right>_{t}\right],$
where $t\mapsto\left<V(\cdot,\widetilde{X}_{\cdot})\right>_{t}$ is the quadratic variation of the process $t\mapsto V(t,\widetilde{X}_{t})$.
6. Conclusion
This paper addresses a fundamental problem in the continuous-time RL framework proposed in [36] by answering the relation between the aggregated dynamics (referred to as the “exploratory dynamics” therein) and the actual sampling process. In terms of SDE theory, we provide a rigorous treatment of the “law-of-large-numbers” heuristics used in [36] to demonstrate how the original dynamics lead to the aggregated dynamics. Our analysis is particularly crucial for understanding how discrete sampling relates to the continuous-time RL problems. We establish the convergence and convergence rate of discretely sampled stochastic policies, i.e., the sampled dynamics, to the aggregated dynamics in different senses. Furthermore, we illustrate the bias and variance of the estimate for several important quantities in analyzing continuous-time RL algorithms when using sampled dynamics, including values, temporal differences, policy gradients, and quadratic variations. These bias and variance analyses are critical for the convergence of stochastic approximation algorithms, and they provide guidance on properly tuning the interaction frequency to foster faster convergence of RL algorithms.
The results of this paper also stimulate more interesting open questions for future research. For example, other sampling methods that could achieve higher accuracy in the RL context, any exact simulation methods for controlled SDEs under stochastic policies if the coefficients are all known/unknown, and a full account for the convergence of the gradient-based continuous-time RL algorithm.
7. Proofs of Main Results
7.1. Proofs of Theorems 4.1, 4.2, 4.3, and 4.4
7.1.1. Proof of Theorem 4.1
Proof.
Without loss of generality, assume $t^{\prime}=t_{j}$, where $1\leq j\leq n$. Consider the following PDE
$\frac{\partial}{\partial t}v(t,x)+\mathcal{L}v(t,x)=0,\quad t\in[0,t_{j}]\times{\mathbb{R}}^{d};\quad v(t_{j},x)=f(x),\quad x\in{\mathbb{R}}^{d}.$ (7.1)
Let $v_{f}(t,x;t_{j})$ be the solution of (7.1). For simplicity, we dorp the $f$ and $t_{j}$ and use the notation $v(t,x)$. By (H. 3), $v(\cdot,\cdot)$ is smooth and $\|v(\cdot,\cdot)\|_{C_{p}^{4}}<+\infty$.
Noting that if $v$ is the solution of (7.1), then by the Feynman-Kac formula, $\mathbb{E}[f(\widetilde{X}_{t_{j}})]=v(0,x_{0})$. With this relation, we have
$\mathbb{E}\left[f(X^{\mathscr{G}}_{t_{j}})-f(\widetilde{X}_{t_{j}})\right]$ $=\mathbb{E}\left[v(t_{j},X^{\mathscr{G}}_{t_{j}})-v(0,x_{0})\right]$ (7.2)
$=\mathbb{E}\left[\sum_{i=1}^{j}\left(v(t_{i},X^{\mathscr{G}}_{t_{i}})-v(t_{i-1},X^{\mathscr{G}}_{t_{i-1}})\right)\right]=:\sum_{i=1}^{j}\mathbb{E}\left[e_{i}\right].$
It suffices to estimate every $\mathbb{E}[e_{i}]$. For simplicity, denote $\partial_{t}:=\partial/\partial t$ and $\partial_{x}:=\partial/\partial x$. It follows from Itô’s lemma (see the dynamic (3.3) of $X^{\mathscr{G}}$) that
$\mathbb{E}\left[e_{i}\right]=$ $\mathbb{E}\left[v(t_{i},X^{\mathscr{G}}_{t_{i}})-v(t_{i-1},X^{\mathscr{G}}_{t_{i-1}})\right]$
$=$ $\mathbb{E}\bigg{[}\int_{t_{i-1}}^{t_{i}}\left(\partial_{t}v(s,X^{\mathscr{G}}_{s})+b^{\top}(s,X^{\mathscr{G}}_{s},a_{t_{i-1}})\partial_{x}v(s,X^{\mathscr{G}}_{s})\right)\mathrm{d}s$
$\quad+\int_{t_{i-1}}^{t_{i}}\frac{1}{2}\operatorname*{Tr}\left(\sigma\sigma^{\top}(s,X^{\mathscr{G}}_{s},a_{t_{i-1}})\partial_{x}^{2}v(s,X^{\mathscr{G}}_{s})\right)\mathrm{d}s\bigg{]},$
with $a_{t_{i-1}}\sim\bm{\pi}(\cdot|t_{i-1},X^{\mathscr{G}}_{t_{i-1}})$, where in the second equality, we use the fact that the Itô integral is a martingale, given that $\sigma$ and $\partial_{x}v$ have polynomial growth and the moments of $X^{\mathscr{G}}_{s}$ are finite (by Lemma 3.1). Noting that $\xi_{i-1}$ is independent of $X^{\mathscr{G}}_{t_{i-1}}$, we have
$\mathbb{E}\bigg{[}\int_{t_{i-1}}^{t_{i}}\left(\partial_{t}v(t_{i-1},X^{\mathscr{G}}_{t_{i-1}})+b^{\top}(t_{i-1},X^{\mathscr{G}}_{t_{i-1}},a_{t_{i-1}})\partial_{x}v(t_{i-1},X^{\mathscr{G}}_{t_{i-1}})\right)\mathrm{d}s$ (7.3)
$\quad+\int_{t_{i-1}}^{t_{i}}\frac{1}{2}\operatorname*{Tr}\left(\sigma\sigma^{\top}(t_{i-1},X^{\mathscr{G}}_{t_{i-1}},a_{t_{i-1}})\partial_{x}^{2}v(t_{i-1},X^{\mathscr{G}}_{t_{i-1}})\right)\mathrm{d}s\bigg{]}$
$=$ $\mathbb{E}\bigg{[}\int_{t_{i-1}}^{t_{i}}\left(\partial_{t}v(t_{i-1},X^{\mathscr{G}}_{t_{i-1}})+(\tilde{b}^{\bm{\pi}})^{\top}(t_{i-1},X^{\mathscr{G}}_{t_{i-1}})\partial_{x}v(t_{i-1},X^{\mathscr{G}}_{t_{i-1}})\right)\mathrm{d}s$
$\quad+\int_{t_{i-1}}^{t_{i}}\frac{1}{2}\operatorname*{Tr}\left((\tilde{\sigma}^{\bm{\pi}})^{2}(t_{i-1},X^{\mathscr{G}}_{t_{i-1}})\partial_{x}^{2}v(t_{i-1},X^{\mathscr{G}}_{t_{i-1}})\right)\mathrm{d}s\bigg{]}=0,$
where the first equality follows form taking conditional expectation $\mathbb{E}[\cdot|X^{\mathscr{G}}_{t_{i-1}}]$ and the second equality follows from (7.1). Therefore,
$\mathbb{E}\left[e_{i}\right]$ $=\mathbb{E}\bigg{[}\int_{t_{i-1}}^{t_{i}}\left(\partial_{t}v(s,X^{\mathscr{G}}_{s})-\partial_{t}v(t_{i-1},X^{\mathscr{G}}_{t_{i-1}})\right)\mathrm{d}s\bigg{]}$ (7.4)
$\quad+\mathbb{E}\bigg{[}\int_{t_{i-1}}^{t_{i}}b^{\top}(s,X^{\mathscr{G}}_{s},a_{t_{i-1}})\left(\partial_{x}v(s,X^{\mathscr{G}}_{s})-\partial_{x}v(t_{i-1},X^{\mathscr{G}}_{t_{i-1}})\right)\mathrm{d}s\bigg{]}$
$\quad+\mathbb{E}\bigg{[}\int_{t_{i-1}}^{t_{i}}\frac{1}{2}\operatorname*{Tr}\left(\sigma\sigma^{\top}(s,X^{\mathscr{G}}_{s},a_{t_{i-1}})\left(\partial_{x}^{2}v(s,X^{\mathscr{G}}_{s})-\partial_{x}^{2}v(t_{i-1},X^{\mathscr{G}}_{t_{i-1}})\right)\right)\mathrm{d}s\bigg{]}$
$\quad+\mathbb{E}\bigg{[}\int_{t_{i-1}}^{t_{i}}\left(b^{\top}(s,X^{\mathscr{G}}_{s},a_{t_{i-1}})-b^{\top}(t_{i-1},X^{\mathscr{G}}_{t_{i-1}},a_{t_{i-1}})\right)\partial_{x}v(t_{i-1},X^{\mathscr{G}}_{t_{i-1}})\mathrm{d}s\bigg{]}$
$\quad+\mathbb{E}\bigg{[}\int_{t_{i-1}}^{t_{i}}\frac{1}{2}\operatorname*{Tr}\left(\left(\sigma\sigma^{\top}(s,X^{\mathscr{G}}_{s},a_{t_{i-1}})-\sigma\sigma^{\top}(t_{i-1},X^{\mathscr{G}}_{t_{i-1}},a_{t_{i-1}})\right)\partial_{x}^{2}v(t_{i-1},X^{\mathscr{G}}_{t_{i-1}})\right)\mathrm{d}s\bigg{]}.$
For $h=\partial_{t}v,\partial_{x}v,\partial_{x}^{2}v,b(\cdot,\cdot,a_{t_{i-1}})\ \text{or}\ \frac{1}{2}\sigma\sigma^{\top}(\cdot,\cdot,a_{t_{i-1}})$, we again use Itô’s lemma to obtain
$h$ $(s,X^{\mathscr{G}}_{s})-h(t_{i-1},X^{\mathscr{G}}_{t_{i-1}})$ (7.5)
$=\int_{t_{i-1}}^{s}\left(\partial_{t}h(u,X^{\mathscr{G}}_{u})+b^{\top}(u,X^{\mathscr{G}}_{u},a_{t_{i-1}})\partial_{x}h(u,X^{\mathscr{G}}_{u})\right)\mathrm{d}u$
$\quad+\int_{t_{i-1}}^{s}\frac{1}{2}\operatorname*{Tr}\left(\sigma\sigma^{\top}(u,X^{\mathscr{G}}_{u},a_{t_{i-1}})\partial_{x}^{2}h(u,X^{\mathscr{G}}_{u})\right)\mathrm{d}u+\int_{t_{i-1}}^{s}(\partial_{x}h(u,X^{\mathscr{G}}_{u}))^{\top}\sigma(u,X^{\mathscr{G}}_{u},a_{t_{i-1}})\mathrm{d}W_{u}.$
By (H.3), Lemma 3.1 and due to $b,\sigma\sigma^{\top}\in C_{p}^{2,0}$, the integrands on the right hand side of the above equation have finite moments. Therefore, if we substitute (7.5) to the first and last two terms of (7.4), then every time integral term is bounded by $C_{p}\|v\|_{C_{p}^{4}}(t_{i}-t_{i-1})^{2}$ for constant $C_{p}>0$ depends only on $b,\sigma,\bm{\pi},T,p$, and the Itô integral terms vanish. For the second and third terms of (7.4), let $h_{1}=b(\cdot,\cdot,a_{t_{i-1}})\ \text{or}\ \frac{1}{2}\sigma\sigma^{\top}(\cdot,\cdot,a_{t_{i-1}})$ and $h_{2}=\partial_{x}v$. We have
$\mathbb{E}\left[\int_{t_{i-1}}^{t_{i}}h_{1}(s,X^{\mathscr{G}}_{s})\left(h_{2}(s,X^{\mathscr{G}}_{s})-h_{2}(t_{i-1},X^{\mathscr{G}}_{t_{i-1}})\right)\mathrm{d}s\right]$ (7.6)
$=$ $\mathbb{E}\left[\int_{t_{i-1}}^{t_{i}}h_{1}(t_{i-1},X^{\mathscr{G}}_{t_{i-1}})\left(h_{2}(s,X^{\mathscr{G}}_{s})-h_{2}(t_{i-1},X^{\mathscr{G}}_{t_{i-1}})\right)\mathrm{d}s\right]$
$+\mathbb{E}\left[\int_{t_{i-1}}^{t_{i}}\left(h_{1}(s,X^{\mathscr{G}}_{s})-h_{1}(t_{i-1},X^{\mathscr{G}}_{t_{i-1}})\right)\left(h_{2}(s,X^{\mathscr{G}}_{s})-h_{2}(t_{i-1},X^{\mathscr{G}}_{t_{i-1}})\right)\mathrm{d}s\right].$
For the first term of (7.6), we again use the Itô’s Lemma to obtain the upper bound $C_{p}\|v\|_{C_{p}^{4}}(t_{i}-t_{i-1})^{2}$. Note that by Hölder’s inequality and (7.5), the second term of (7.6) is bounded by
$\sqrt{\mathbb{E}\left[\int_{t_{i-1}}^{t_{i}}\left|h_{1}(s,X^{\mathscr{G}}_{s})-h_{1}(t_{i-1},X^{\mathscr{G}}_{t_{i-1}})\right|^{2}\mathrm{d}s\right]\mathbb{E}\left[\int_{t_{i-1}}^{t_{i}}\left|h_{2}(s,X^{\mathscr{G}}_{s})-h_{2}(t_{i-1},X^{\mathscr{G}}_{t_{i-1}})\right|^{2}\mathrm{d}s\right]}\leq C_{p}\|v\|_{C_{p}^{4}}(t_{i}-t_{i-1})^{2}.$
Thus $\mathbb{E}\left[e_{i}\right]\leq C_{p}\|v\|_{C_{p}^{4}}(t_{i}-t_{i-1})^{2},$ where $C_{p}$ is a constant depending only on $b,\sigma,\bm{\pi},T,p$. The desired result follows from $\sum_{i=1}^{j}(t_{i}-t_{i-1})^{2}\leq T|\mathscr{G}|$. ∎
7.1.2. Proof of Theorem 4.2
The following lemma is similar to [23, Lemma 14.1.6]. One can combine the method in Theorem 4.1 to prove it, so we omit the proof here.
Lemma 7.1.
Suppose (H.1) holds, $l\in(0,1)\cup(1,2)\cup(2,3)$ and $g\in\mathcal{H}_{T}^{(l)}$. Then there exist a constant $C$, such that for all $s\in[0,T]$,
$\left|\mathbb{E}\left[g(s,X^{\mathscr{G}}_{s},a_{\delta(s)})-g(\delta(s),X^{\mathscr{G}}_{\delta(s)},a_{\delta(s)})\right]\right|\leq C\|g\|_{T}^{(l)}|\mathscr{G}|^{\mathcal{X}(l)},$ (7.7)
where $\mathcal{X}(l)$ satisfies (4.13).
Based on Lemma 7.1, we prove Theorem 4.2.
Proof.
It suffices to prove $|\mathbb{E}[f(X^{\mathscr{G}}_{t^{\prime}})]-\mathbb{E}[f(\widetilde{X}_{t^{\prime}})]|\leq C\|f\|^{(l+2)}|\mathscr{G}|^{\mathcal{X}(l)}$ for any $t^{\prime}\in[0,T]$. Under (H. 4), by [24, Chapter 4, Theorem 5.1], for any $f\in\mathcal{H}^{(l+2)}$, the PDE (4.4) has a unique solution $v(\cdot,\cdot;t^{\prime})\in\mathcal{H}_{t^{\prime}}^{(l+2)}$ and $\|v(\cdot,\cdot;t^{\prime})\|_{t^{\prime}}^{(l+2)}\leq C\|f\|^{(l+2)}$, where $C$ is a constant depending only on $T,\tilde{b}^{\bm{\pi}},\tilde{\sigma}^{\bm{\pi}}$. Note that $\mathcal{H}_{T}^{(l)}\subset\mathcal{H}_{s}^{(l)}$ for every $0\leq t^{\prime}\leq T$. Without loss of generality, suppose that $t^{\prime}=T$. For simplicity, we drop the notation $t^{\prime}$ and use $v$ to represent the solution of (4.4) with $t^{\prime}=T$. Recall (7.2),
$\mathbb{E}\left[f(X^{\mathscr{G}}_{T})-f(\widetilde{X}_{T})\right]$ $=\mathbb{E}\left[v(T,X^{\mathscr{G}}_{T})-v(0,x_{0})\right]=\sum_{i=1}^{n}\mathbb{E}\left[e_{i}\right].$
Note that (7.3) still holds in this situation. Therefore, we have
$\mathbb{E}$ $\left[e_{i}\right]=\mathbb{E}\bigg{[}\int_{t_{i-1}}^{t_{i}}\left(\partial_{t}v(s,X^{\mathscr{G}}_{s})-\partial_{t}v(t_{i-1},X^{\mathscr{G}}_{t_{i-1}})\right)\mathrm{d}s\bigg{]}$ (7.8)
$+\mathbb{E}\bigg{[}\int_{t_{i-1}}^{t_{i}}b^{\top}(s,X^{\mathscr{G}}_{s},a_{t_{i-1}})\partial_{x}v(s,X^{\mathscr{G}}_{s})-b^{\top}(t_{i-1},X^{\mathscr{G}}_{t_{i-1}},a_{t_{i-1}})\partial_{x}v(t_{i-1},X^{\mathscr{G}}_{t_{i-1}})\mathrm{d}s\bigg{]}$
$+\mathbb{E}\bigg{[}\int_{t_{i-1}}^{t_{i}}\frac{1}{2}\operatorname*{Tr}\left(\sigma\sigma^{\top}(s,X^{\mathscr{G}}_{s},a_{t_{i-1}})\partial_{x}^{2}v(s,X^{\mathscr{G}}_{s})-\sigma\sigma^{\top}(t_{i-1},X^{\mathscr{G}}_{t_{i-1}},a_{t_{i-1}})\partial_{x}^{2}v(t_{i-1},X^{\mathscr{G}}_{t_{i-1}})\right)\mathrm{d}s\bigg{]}.$
Note that $\partial_{t}v,b^{\top}\partial_{x}v$ and $\frac{1}{2}\operatorname*{Tr}(\sigma\sigma^{\top}\partial_{x}^{2}v)$ belong to $\mathcal{H}_{T}^{(l)}$. By Lemma 7.1, the three terms on the right hand side of above equation are all bounded by $C\|v\|^{(l+2)}_{T}|\mathscr{G}|^{\mathcal{X}(l)}(t_{i}-t_{i-1})$ with a constant $C$ depending only on $b,\sigma,T$. As a result, we have
$\bigg{|}\mathbb{E}\left[f(X^{\mathscr{G}}_{T})-f(\widetilde{X}_{T})\right]\bigg{|}\leq\sum_{i=1}^{n}\left|\mathbb{E}\left[e_{i}\right]\right|\leq\sum_{i=1}^{n}C\|v\|^{(l+2)}_{T}|\mathscr{G}|^{\mathcal{X}(l)}(t_{i}-t_{i-1})\leq CT\|f\|^{(l+2)}|\mathscr{G}|^{\mathcal{X}(l)}.$ (7.9)
This proves the desired result. ∎
7.1.3. Proof of Theorem 4.3
Recall $\mathcal{F}^{\xi}:=\sigma\{\xi_{i}\mid{i\geq 0}\}$ as defined in Section 4.2. To facilitate the proof, we define $\mathcal{F}^{\xi}_{i}:=\sigma\{\xi_{j}\mid j\leq i-1\}$ for all $i\geq 0$, and define the filtration {~$\{\widetilde{\mathcal{F}}_{t}\}_{t\geq 0}$ with ~$\widetilde{\mathcal{F}}_{t}:=\mathcal{F}_{t}^{W}\vee\mathcal{F}^{\xi}$. Note that $\{W_{s}\}_{s\geq 0}$ is still a Brownian motion on $(\Omega,\mathcal{F},{\mathbb{P}})$ with filtration {~$\{\widetilde{\mathcal{F}}_{t}\}_{t\geq 0}$ and we have $\mathbb{E}^{W}\left[\cdot\right]=\mathbb{E}\big{[}\cdot\big{|}\widetilde{\mathcal{F}}_{0}\big{]}$.
The following lemma shows that $X^{\mathscr{G}}$ has finite moments conditional on $\mathcal{F}^{\xi}$. The proof is given in Appendix A.
Lemma 7.2.
Under condition (4.14) of (H.5), for $p\geq 2$, there is a constant $C_{p}$ such that
$\mathbb{E}^{W}\left[|X^{\mathscr{G}}_{s}|^{p}\right]\leq C_{p}(1+|x_{0}|^{p})e^{C_{p}s},\quad s\in[0,T].$ (7.10)
Based on Lemma 7.2, we prove Theorem 4.3.
Proof of Theorem 4.3.
Without loss of generality, assume that $t=t_{j}$ for some $1\leq j\leq n$. To prove (4.15), consider the following PDE:
$\frac{\partial}{\partial t}v(t,x)+\mathcal{L}v(t,x)=0,\quad(t,x)\in[0,t_{j}]\times{\mathbb{R}}^{d};\quad v(t_{j},x)=f(x),\quad x\in{\mathbb{R}}^{d}.$ (7.11)
Under the conditions of this theorem, the solution $v$ of (7.11) is smooth and bounded. By the Feynman-Kac formula, $\mathbb{E}^{W}[f(\widetilde{X}_{t_{j}})]=v(0,x_{0})$. With this relation, we have
$\mathbb{E}^{W}\left[f(X^{\mathscr{G}}_{t_{j}})-f(\widetilde{X}_{t_{j}})\right]$ $=\mathbb{E}^{W}\left[v(t_{j},X^{\mathscr{G}}_{t_{j}})-v(0,x_{0})\right]$
$=\mathbb{E}^{W}\left[\sum_{i=1}^{j}\left(v(t_{i},X^{\mathscr{G}}_{t_{i}})-v(t_{i-1},X^{\mathscr{G}}_{t_{i-1}})\right)\right]=:\sum_{i=1}^{j}\mathbb{E}^{W}\left[e_{i}\right].$
As in the proof of Theorem 4.1, noting that the Itô integral in (7.5) is mean zero under $\mathbb{P}^{W}$, we can decompose $\mathbb{E}^{W}\left[e_{i}\right]$ into two terms and obtain
$\mathbb{E}^{W}\left[f(X^{\mathscr{G}}_{T})\right]-\mathbb{E}^{W}\left[f(\widetilde{X}_{T})\right]=\sum_{i=1}^{j}A_{i}+\sum_{i=1}^{j}B_{i},$ (7.12)
where $A_{i}$ is defined by
$A_{i}:=$ $\mathbb{E}^{W}\bigg{[}\int_{t_{i-1}}^{t_{i}}\left(\partial_{t}v(t_{i-1},X^{\mathscr{G}}_{t_{i-1}})+b^{\top}(t_{i-1},X^{\mathscr{G}}_{t_{i-1}},a_{t_{i-1}})\partial_{x}v(t_{i-1},X^{\mathscr{G}}_{t_{i-1}})\right)\mathrm{d}s$
$\quad+\int_{t_{i-1}}^{t_{i}}\frac{1}{2}\operatorname*{Tr}\left(\sigma\sigma^{\top}(t_{i-1},X^{\mathscr{G}}_{t_{i-1}},a_{t_{i-1}})\partial_{x}^{2}v(t_{i-1},X^{\mathscr{G}}_{t_{i-1}})\right)\mathrm{d}s\bigg{]}.$
and $B_{i}$ is defined by
$B_{i}$ $\coloneqq\mathbb{E}^{W}\bigg{[}\int_{t_{i-1}}^{t_{i}}\left(\partial_{t}v(s,X^{\mathscr{G}}_{s})-\partial_{t}v(t_{i-1},X^{\mathscr{G}}_{t_{i-1}})\right)\mathrm{d}s\bigg{]}$ (7.13)
$\quad+\mathbb{E}^{W}\bigg{[}\int_{t_{i-1}}^{t_{i}}b^{\top}(s,X^{\mathscr{G}}_{s},a_{t_{i-1}})\left(\partial_{x}v(s,X^{\mathscr{G}}_{s})-\partial_{x}v(t_{i-1},X^{\mathscr{G}}_{t_{i-1}})\right)\mathrm{d}s\bigg{]}$
$\quad+\mathbb{E}^{W}\bigg{[}\int_{t_{i-1}}^{t_{i}}\frac{1}{2}\operatorname*{Tr}\left(\sigma\sigma^{\top}(s,X^{\mathscr{G}}_{s},a_{t_{i-1}})\left(\partial_{x}^{2}v(s,X^{\mathscr{G}}_{s})-\partial_{x}^{2}v(t_{i-1},X^{\mathscr{G}}_{t_{i-1}})\right)\right)\mathrm{d}s\bigg{]}$
$\quad+\mathbb{E}^{W}\bigg{[}\int_{t_{i-1}}^{t_{i}}\left(b^{\top}(s,X^{\mathscr{G}}_{s},a_{t_{i-1}})-b^{\top}(t_{i-1},X^{\mathscr{G}}_{t_{i-1}},a_{t_{i-1}})\right)\partial_{x}v(t_{i-1},X^{\mathscr{G}}_{t_{i-1}})\mathrm{d}s\bigg{]}$
$\quad+\mathbb{E}^{W}\bigg{[}\int_{t_{i-1}}^{t_{i}}\frac{1}{2}\operatorname*{Tr}\left(\left(\sigma\sigma^{\top}(s,X^{\mathscr{G}}_{s},a_{t_{i-1}})-\sigma\sigma^{\top}(t_{i-1},X^{\mathscr{G}}_{t_{i-1}},a_{t_{i-1}})\right)\partial_{x}^{2}v(t_{i-1},X^{\mathscr{G}}_{t_{i-1}})\right)\mathrm{d}s\bigg{]}.$
We again use the Itô’s lemma to analysis every difference term of (7.13). Note that every term in $B_{j}$ can be written as double integrals with respect to time as in the proof of Theorem 4.1, and by conditions (H.5) and (H.3), the integand is bounded by $C_{p}\|v\|_{C_{p}^{4}}(1+|X^{\mathscr{G}}_{s}|^{m_{p}})$ for some $m_{p}$ and $C_{p}$ that depend on $p$. Applying Lemma 7.2, we obtain
$|B_{i}|\leq C_{p}\|v\|_{C_{p}^{4}}(t_{i}-t_{i-1})^{2},$
where $C_{p}$ is a constant depending on $b,\sigma,\bm{\pi},T,p$. As a result,
$\sum_{i=1}^{j}|B_{i}|\leq\sum_{i=1}^{j}C_{P}\|v\|_{C_{p}^{4}}|\mathscr{G}|(t_{i}-t_{i-1})\leq TC_{p}\|v\|_{C_{p}^{4}}|\mathscr{G}|.$ (7.14)
As for $A_{i}$, note that $\{(A_{i},\mathcal{F}^{\xi}_{i})\}_{i}$ is a martingale difference sequence. In fact, denote
$S_{i}:=\bigg{(}\partial_{t}v(t_{i-1},X^{\mathscr{G}}_{t_{i-1}})$ $+b^{\top}(t_{i-1},X^{\mathscr{G}}_{t_{i-1}},a_{t_{i-1}})\partial_{x}v(t_{i-1},X^{\mathscr{G}}_{t_{i-1}})$ (7.15)
$+\frac{1}{2}\operatorname*{Tr}\left(\sigma\sigma^{\top}(t_{i-1},X^{\mathscr{G}}_{t_{i-1}},a_{t_{i-1}})\partial_{x}^{2}v(t_{i-1},X^{\mathscr{G}}_{t_{i-1}})\right)\bigg{)}(t_{i}-t_{i-1}).$
Observing that $S_{i}$ is independent of $\xi_{i},\xi_{i+1},\dots$, we have $A_{i}=\mathbb{E}^{W}\left[S_{i}\right]=\mathbb{E}[S_{i}\mid\mathcal{F}^{\xi}_{i}]$ is $\mathcal{F}^{\xi}_{i}$-measurable and for $i\geq 2$,
$\mathbb{E}[A_{i}\mid\mathcal{F}^{\xi}_{i-1}]$ $=\mathbb{E}\left[\mathbb{E}^{W}\left[S_{i}\right]\big{|}\xi_{0},\dots,\xi_{i-2}\right]=\mathbb{E}\left[\mathbb{E}\left[S_{i}\big{|}\xi_{0},\dots,\xi_{i-2},\xi_{i-1}\right]\big{|}\xi_{0},\dots,\xi_{i-2}\right]$ (7.16)
$=\mathbb{E}\left[\mathbb{E}\left[S_{i}\big{|}X_{t_{i-1}},\xi_{0},\dots,\xi_{i-2}\right]\big{|}\xi_{0},\dots,\xi_{i-2}\right]=0,$
where the last equality follows from the same result in (7.3).
By conditions (H.3) and (H.5), there is a constant $C_{p}$ and an integer $m_{p}$ such that $|S_{i}|\leq C_{p}\|v\|_{C_{p}^{4}}(1+|X^{\mathscr{G}}_{t_{i-1}}|^{m_{p}})$. Again using Lemma 7.2, we obtain
$|A_{i}|\leq C_{p}\|v\|_{C_{p}^{4}}(t_{i}-t_{i-1}).$ (7.17)
By [35, Corollary 2.20], we have
$\mathbb{P}\bigg{(}\bigg{|}\sum_{i=1}^{j}A_{i}\bigg{|}\geq t\bigg{)}\leq 2\exp\left(-\frac{t^{2}}{4T(C_{p}\|v\|_{C_{p}^{4}})^{2}|\mathscr{G}|}\right).$ (7.18)
Let $\varrho=2\exp\left(-\frac{t^{2}}{4T(C_{p}\|v\|_{C_{p}^{4}})^{2}|\mathscr{G}|}\right)$, we obtain
$\mathbb{P}\bigg{(}\bigg{|}\sum_{j=1}^{i}A_{i}\bigg{|}\geq 2\sqrt{T}C_{p}\|v\|_{C_{p}^{4}}\sqrt{|\mathscr{G}|}\sqrt{\ln(2/\varrho)}\bigg{)}\leq\varrho.$ (7.19)
It follows from (7.14) that
$\mathbb{P}$ $\left(\left|\mathbb{E}^{W}[f(X^{\mathscr{G}}_{t^{\prime}})]-\mathbb{E}^{W}[f(\widetilde{X}_{t^{\prime}})]\right|>(2+\sqrt{T})\sqrt{T}C_{p}\|v\|_{C_{p}^{4}}\sqrt{|\mathscr{G}|}\big{(}1+\sqrt{\ln(2/\varrho)}\big{)}\right)$
$\quad\leq\mathbb{P}\bigg{(}\bigg{|}\sum_{j=1}^{i}A_{i}\bigg{|}\geq 2\sqrt{T}C_{p}\|v\|_{C_{p}^{4}}\sqrt{|\mathscr{G}|}\sqrt{\ln(2/\varrho)}\bigg{)}+\mathbb{P}\left(\sum_{i=1}^{j}|B_{i}|\geq(2+\sqrt{T})\sqrt{T}C_{p}\|v\|_{C_{p}^{4}}\sqrt{|\mathscr{G}|}\right)$
$\quad\leq\varrho+0=\varrho.$
The desired result follows from $\|v\|_{C_{p}^{4}}\leq C\|f\|_{C_{p}^{4}}$.
For the proof of (4.16), we can start from (7.12). Note that $\{A_{i}\}_{i}$ is still a martingale difference sequence. By (H.4), we have
$|A_{i}|\leq C\|v\|_{t_{j}}^{(l+2)}(t_{i}-t_{i-1})$ (7.20)
It suffices to analyze (7.13). Applying Lemmas 7.2 and 7.1, one can use a similar method in the proof of Theorem 4.2 to obtain
$|B_{i}|\leq C\|v\|^{(l+2)}_{t_{j}}|\mathscr{G}|^{\mathcal{X}(l)}(t_{i}-t_{i-1}).$ (7.21)
As a result,
$\sum_{i=1}^{j}|B_{i}|\leq CT\|v\|^{(l+2)}_{t_{j}}|\mathscr{G}|^{\mathcal{X}(l)}.$ (7.22)
By [35, Corollary 2.20], we have
$\mathbb{P}\bigg{(}\bigg{|}\sum_{i=1}^{j}A_{i}\bigg{|}\geq t\bigg{)}\leq 2\exp\left({-\frac{t^{2}}{4T(C\|v\|_{t_{j}}^{(l+2)})^{2}|\mathscr{G}|}}\right).$ (7.23)
Let $\varrho=2\exp\left({-\frac{t^{2}}{4T(C\|v\|_{t_{j}}^{(l+2)})^{2}|\mathscr{G}|}}\right)$, we obtain
$\mathbb{P}\bigg{(}\bigg{|}\sum_{i=1}^{j}A_{i}\bigg{|}\geq 2\sqrt{T}C\|v\|_{t_{j}}^{(l+2)}\sqrt{|\mathscr{G}|}\sqrt{\ln(2/\varrho)}\bigg{)}\leq\varrho.$ (7.24)
It follows from (7.22) that
$\mathbb{P}$ $\left(\left|\mathbb{E}^{W}[f(X^{\mathscr{G}}_{t^{\prime}})]-\mathbb{E}^{W}[f(\widetilde{X}_{t^{\prime}})]\right|>(2+\sqrt{T})\sqrt{T}C\|v\|_{t_{j}}^{(l+2)}|\mathscr{G}|^{\mathcal{X}(l)\land\frac{1}{2}}\big{(}1+\sqrt{\ln(2/\varrho)}\big{)}\right)$
$\quad\leq\mathbb{P}\bigg{(}\bigg{|}\sum_{i=1}^{j}A_{i}\bigg{|}\geq 2\sqrt{T}C\|v\|_{t_{j}}^{(l+2)}\sqrt{|\mathscr{G}|}\sqrt{\ln(2/\varrho)}\bigg{)}+\mathbb{P}\left(\sum_{i=1}^{j}|B_{i}|\geq(2+\sqrt{T})\sqrt{T}C\|v\|_{t_{j}}^{(l+2)}|\mathscr{G}|^{\mathcal{X}(l)}\right)$
$\quad\leq\varrho+0=\varrho.$
The desired result follows from $\|v\|_{t_{j}}^{(l+2)}\leq C\|f\|^{(l+2)}$. ∎
7.1.4. Proof of Theorem 4.4
We introduce the Orlicz norm on $(\Omega^{\xi},\mathcal{F}^{\xi},\mathbb{P}^{\xi})$. Let $\psi$ be a convex, nondecreasing function with $\psi(0)=0$. For a random variable $X(\omega):\Omega^{\xi}\rightarrow\mathbb{R}^{m},m\in\mathbb{N}_{+}$, the Orlicz norm is defined as
$\|X\|_{\psi}=\inf\left\{C>0:\mathbb{E}\left[\psi\left(\frac{|X|}{C}\right)\right]\leq 1\right\},$ (7.25)
where ||$\cdot$|| ⋅ | is the Euclid norm on $\mathbb{R}^{m}$, and the corresponding Orlicz space is $L_{\psi}$. Note that the $L_{p}$-norm $\|\cdot\|_{p}$ is the corresponding Orlicz norm with the function $\psi(x)=x^{p}$. Denote $\psi_{\alpha}=e^{x^{\alpha}}-1,\alpha>0$ (linear near the origin when $10<\alpha<$ 1 in order for $\psi_{\alpha}$ to be convex). We say $X$ is sub-exponential if and only if $X\in L_{\psi_{1}}$. It follows from $|x|^{p}/p!\leq e^{|x|}-1$ that
$\|X\|_{p}\leq\sqrt[p]{p!}\|X\|_{\psi_{1}}.$ (7.26)
Combining the above inequality and [34, Lemma 2.2.2], we derive the following lemma.
Lemma 7.3.
For $p\geq 1$, if $\{X_{i}\}_{i=1}^{n}$ are sub-exponential random variables, then there is a constant $C_{p}$, such that
$\mathbb{E}\left[\max_{1\leq i\leq n}|X_{i}|^{p}\right]\leq C_{p}\log^{p}(1+n)\max_{i}\|X_{i}\|_{\psi_{1}}^{p}.$ (7.27)
Note that under condition (H.6), $\tilde{b}^{\bm{\pi}}$ and $\tilde{\sigma}^{\bm{\pi}}$ is globally Lipschitz with respect to the state $x$. By classical SDE results, e.g. [13, Theorem 7.9], we have the following lemma.
Lemma 7.4.
Assume (H.6) holds. For every $p\geq 2$, there is a constant $C_{p}$, such that
$\mathbb{E}^{W}\left[\sup_{0\leq t\leq T}|\widetilde{X}_{t}|^{p}\right]\leq C_{p},\quad\mathbb{E}\left[\sup_{0\leq t\leq T}|X^{\mathscr{G}}_{t}|^{p}\right]\leq C_{p}.$ (7.28)
Moreover,
$\mathbb{E}^{W}\left[\sup_{0\leq s\leq T}|\widetilde{X}_{s}-\widetilde{X}_{\delta(s)}|^{p}\right]\leq C_{p}|\mathscr{G}|^{\frac{p}{2}},\quad\mathbb{E}\left[\sup_{0\leq s\leq T}|X^{\mathscr{G}}_{s}-X^{\mathscr{G}}_{\delta(s)}|^{p}\right]\leq C_{p}|\mathscr{G}|^{\frac{p}{2}}.$ (7.29)
Proof of Theorem 4.4.
By the Lipschitz condition of $\phi$, $d_{A}(\phi(t_{0},x_{0},\xi),a_{0})$ is sub-exponential if and only if $d_{A}(\phi(0,0,\xi),a_{0})$ is sub-exponential by the triangular inequality and the condition (1) of (H.6). we take $(t_{0},x_{0})=(0,0)$ for simplicity. Recall the notation $i(t):=\max\{i:t_{i}\leq t\}$. Note that
$X^{\mathscr{G}}_{t}$ $=x_{0}+\int_{0}^{t}b(s,X^{\mathscr{G}}_{s},\phi(\delta(s),X^{\mathscr{G}}_{\delta(s)},\xi_{i(s)}))\mathrm{d}s+\int_{0}^{t}\sigma(s,X^{\mathscr{G}}_{s})\mathrm{d}W_{s},$ (7.30)
$\widetilde{X}_{t}$ $=x_{0}+\int_{0}^{t}\tilde{b}^{\bm{\pi}}(s,\widetilde{X}_{s})\mathrm{d}s+\int_{0}^{t}\sigma(s,\widetilde{X}_{s})\mathrm{d}W_{s}.$
Denote $\Delta X_{t}:=X^{\mathscr{G}}_{t}-\widetilde{X}_{t}$. We have
$\sup_{0\leq u\leq t}|\Delta X_{u}|$ $=\sup_{0\leq u\leq t}\bigg{|}\int_{0}^{u}\left(b(s,X^{\mathscr{G}}_{s},\phi(\delta(s),X^{\mathscr{G}}_{\delta(s)},\xi_{i(s)}))-b(s,\widetilde{X}_{s},\phi(s,\widetilde{X}_{s},\xi_{i(s)}))\right)\mathrm{d}s$ (7.31)
$\quad+\int_{0}^{u}\left(b(s,\widetilde{X}_{s},\phi(s,\widetilde{X}_{s},\xi_{i(s)}))-\tilde{b}^{\bm{\pi}}(s,\widetilde{X}_{s})\right)\mathrm{d}s+\int_{0}^{u}\left(\sigma(s,X^{\mathscr{G}}_{s})-\sigma(s,\widetilde{X}_{s})\right)\mathrm{d}W_{s}\bigg{|}.$
It follows from the Lipschitz conditions and the Burkholder-Davis-Gundy inequality that for $p\geq 2$,
$\mathbb{E}\left[\sup_{0\leq u\leq t}\left|\int_{0}^{u}\left(\sigma(s,X^{\mathscr{G}}_{s})-\sigma(s,\widetilde{X}_{s})\right)\mathrm{d}W_{s}\right|^{p}\right]$ $\leq C_{p}\mathbb{E}\left[\left(\int_{0}^{t}\left|\sigma(s,X^{\mathscr{G}}_{s})-\sigma(s,\widetilde{X}_{s})\right|^{2}\mathrm{d}s\right)^{\frac{p}{2}}\right]$
$\leq C_{p}\mathbb{E}\left[\left(\int_{0}^{t}|\Delta X_{u}|^{2}\mathrm{d}u\right)^{\frac{p}{2}}\right]\leq C_{p}\mathbb{E}\left[\int_{0}^{t}|\Delta X_{u}|^{p}\mathrm{d}u\right].$
Thus, we have
$\mathbb{E}\left[\sup_{0\leq s\leq t}\left|\Delta X_{s}\right|^{p}\right]\leq$ $C_{p}|\mathscr{G}|^{\frac{p}{2}}+C_{p}\int_{0}^{t}\mathbb{E}\left[\left|\Delta X_{s}\right|^{p}+\left|X^{\mathscr{G}}_{s}-X^{\mathscr{G}}_{\delta(s)}\right|^{p}\right]\mathrm{d}s$ (7.32)
$+C_{p}\mathbb{E}\left[\sup_{0\leq u\leq t}\left|\int_{0}^{u}b(s,\widetilde{X}_{s},\phi(s,\widetilde{X}_{s},\xi_{i(s)}))\mathrm{d}s-\int_{0}^{u}\tilde{b}^{\bm{\pi}}(s,\widetilde{X}_{s})\mathrm{d}s\right|^{p}\right].$
We focus on the third term of the above equation. Denote $U_{s}:=b(s,\widetilde{X}_{s},\phi(s,\widetilde{X}_{s},\xi_{i(s)}))-\tilde{b}^{\bm{\pi}}(s,\widetilde{X}_{s})$. Under (H.6), we have
$|U_{s}|=\left|b(s,\widetilde{X}_{s},\phi(s,\widetilde{X}_{s},\xi_{i(s)}))-\tilde{b}^{\bm{\pi}}(s,\widetilde{X}_{s})\right|\leq C\left(1+\sup_{0\leq t\leq T}|\widetilde{X}_{t}|+d_{A}(\phi(0,0,\xi_{i(s)}),a_{0})\right).$ (7.33)
Denote $\mathbb{E}_{\xi}\coloneqq\mathbb{E}\left[\cdot|\mathcal{F}^{W}\right]$ and note that $\xi$ is independent of $\mathcal{F}^{W}$.
$\mathbb{E}\left[\sup_{0\leq u\leq t}\left|\int_{0}^{u}U_{s}\mathrm{d}s\right|^{p}\right]\leq C_{p}\mathbb{E}\left[\sup_{0\leq u\leq t}\left|\int_{0}^{\delta(u)}U_{s}\mathrm{d}s\right|^{p}\right]+C_{p}\mathbb{E}\left[\mathbb{E}_{\xi}\left[\sup_{0\leq u\leq t}\left|\int_{\delta(u)}^{u}U_{s}\mathrm{d}s\right|^{p}\right]\right].$ (7.34)
It follows from (7.33) and Lemma 7.3 that
$\mathbb{E}\left[\mathbb{E}_{\xi}\left[\sup_{0\leq u\leq t}\left|\int_{\delta(u)}^{u}U_{s}\mathrm{d}s\right|^{p}\right]\right]$ $\leq C_{p}\mathbb{E}\left[\left(1+\sup_{0\leq t\leq T}\left|\widetilde{X}_{t}\right|^{p}+\mathbb{E}\left[\max_{1\leq j\leq n}\left(d_{A}(\phi(0,0,\xi_{j}),a_{0})\right)^{p}\right]\right)|\mathscr{G}|^{p}\right]$ (7.35)
$\leq C_{p}\mathbb{E}\left[\left(1+\sup_{0\leq t\leq T}\left|\widetilde{X}_{t}\right|^{p}+\log^{p}(1+n)\|d_{A}(\phi(0,0,\xi),a_{0})\|_{\psi_{1}}^{p}\right)|\mathscr{G}|^{p}\right]$
$\leq C_{p}|\mathscr{G}|^{\frac{p}{2}},$
where in the last inequality, we used the fact that $|\mathscr{G}|^{1/2}\log(1+n)\leq C_{G}$.
As for the first term on the right hand side of (7.34), denote $M_{j}:=\int_{t_{j}}^{t_{j+1}}U_{s}\mathrm{d}s$ and $\mathcal{F}_{i}:=\sigma(\xi_{j}:0\leq j\leq i-1)\lor\mathcal{F}^{W}$. Note that $\left\{(\sum_{j=0}^{i-1}M_{j},\mathcal{F}_{i})\right\}_{i=1}^{n}$ is a martingale. It follows from Doob’s inequality that
$\mathbb{E}\left[\sup_{0\leq u\leq t}\left|\int_{0}^{\delta(u)}U_{s}\mathrm{d}s\right|^{p}\right]=\mathbb{E}\left[\max_{1\leq i\leq n}\Big{|}\sum_{j=0}^{i-1}M_{j}\Big{|}^{p}\right]\leq C_{p}\mathbb{E}\left[\Big{|}\sum_{j=0}^{n-1}M_{j}\Big{|}^{p}\right],$ (7.36)
where $C_{p}$ is a constant depend on $p$.
By (H.6) and relation (7.33), $M_{j}$ is sub-exponential conditional on $\mathcal{F}^{W}$ and with a little abuse of notations,
$\|M_{j}\|_{\psi_{1}}\leq C\left(1+\sup_{0\leq t\leq T}|\widetilde{X}_{t}|\|1\|_{\psi_{1}}+\|d_{A}(\phi(0,0,\xi),a_{0})\|_{\psi_{1}}\right)(t_{j+1}-t_{j}),$ (7.37)
where $\|\cdot\|_{\psi_{1}}$ is the Orlicz norm on $(\Omega^{\xi},\mathcal{F}^{\xi},\mathbb{P}^{\xi})$. It follows from [25, Theorem 6.21] that
$\Big{\|}\sum_{j=0}^{n-1}M_{j}\Big{\|}_{\psi_{1}}\leq C\left(\mathbb{E}_{\xi}\left[\Big{|}\sum_{j=0}^{n-1}M_{j}\Big{|}\right]+\|\max_{j}|M_{j}|\|_{\psi_{1}}\right).$ (7.38)
For the first term on the right hand side of the above equation, using the fact that $\mathbb{E}_{\xi}\left[M_{i}^{\top}M_{j}\right]=0,i\neq j$, it follows from relation (7.33) and $\psi(0,0,\xi)$ is square integrable that
$\mathbb{E}_{\xi}\bigg{[}\Big{|}\sum_{j=0}^{n-1}M_{j}\Big{|}\bigg{]}\leq\sqrt{\mathbb{E}_{\xi}\bigg{[}\Big{|}\sum_{j=0}^{n-1}M_{j}\Big{|}^{2}\bigg{]}}=\sqrt{\sum_{j=0}^{n-1}\mathbb{E}_{\xi}\left[|M_{j}|^{2}\right]}\leq C(1+\sup_{0\leq t\leq T}|\widetilde{X}_{t}|)|\mathscr{G}|^{\frac{1}{2}}.$ (7.39)
As for the second term, it follows from [34, Lemma 2.2.2] and (7.37) that
$\|\max_{j}|M_{j}|\|_{\psi_{1}}\leq C\log(1+n)\max_{j}\|M_{j}\|_{\psi_{1}}\leq C\log(1+n)\left(1+\sup_{0\leq t\leq T}|\widetilde{X}_{t}|\right)|\mathscr{G}|.$ (7.40)
In conclusion, we have
$\Big{\|}\sum_{j=0}^{n-1}M_{j}\Big{\|}_{\psi_{1}}\leq C(1+\sup_{0\leq t\leq T}|\widetilde{X}_{t}|)|\mathscr{G}|^{\frac{1}{2}}.$ (7.41)
Conditional on $\mathcal{F}^{W}$, we have
$\mathbb{P}_{\xi}\left(\bigg{|}\sum_{j=0}^{n-1}M_{j}\bigg{|}^{p}\geq t\right)$ $=\mathbb{P}_{\xi}\left(\psi_{1}\bigg{(}\bigg{|}\sum_{j=0}^{n-1}M_{j}\bigg{|}/\Big{\|}\sum_{j=0}^{n-1}M_{j}\Big{\|}_{\psi_{1}}\bigg{)}\leq\psi_{1}\bigg{(}t^{\frac{1}{p}}/\Big{\|}\sum_{j=0}^{n-1}M_{j}\Big{\|}_{\psi_{1}}\bigg{)}\right)$ (7.42)
$\leq 1/\psi_{1}\bigg{(}t^{\frac{1}{p}}/\Big{\|}\sum_{j=0}^{n-1}M_{j}\Big{\|}_{\psi_{1}}\bigg{)}\leq 2\exp\left\{-\frac{t^{\frac{1}{p}}}{C(1+\sup_{0\leq t\leq T}|\widetilde{X}_{t}|)|\mathscr{G}|^{\frac{1}{2}}}\right\}.$
Integrating the above equation with respect to $t$ from 00 to +$\infty$+ ∞, we have
$\mathbb{E}\left[\left|\sum_{j=0}^{n-1}\int_{t_{j}}^{t_{j+1}}b(s,\widetilde{X}_{s},\phi(s,\widetilde{X}_{s},\xi_{j}))\mathrm{d}s-\int_{t_{i}}^{t_{i+1}}\tilde{b}^{\bm{\pi}}(s,\widetilde{X}_{s})\mathrm{d}s\right|^{p}\bigg{|}\mathcal{F}^{W}\right]\leq C_{p}(1+\sup_{0\leq t\leq T}|\widetilde{X}_{t}|^{p})|\mathscr{G}|^{\frac{p}{2}}.$ (7.43)
Combining (7.32), (7.34), (7.35) and (7.43), using Lemma 7.4 to bound the time difference of $X^{\mathscr{G}}$ and $\widetilde{X}$ and the moments of $\sup_{t}|\widetilde{X}_{t}|$, we have
$\mathbb{E}\left[\sup_{0\leq s\leq t}\left|\Delta X_{s}\right|^{p}\right]$ $\leq C_{p}|\mathscr{G}|^{\frac{p}{2}}+C_{p}\int_{0}^{t}\mathbb{E}\left[\sup_{0\leq u\leq s}\left|\Delta X_{u}\right|^{p}\right]+\mathbb{E}\left[|X^{\mathscr{G}}_{s}-X^{\mathscr{G}}_{\delta(s)}|^{p}\right]\mathrm{d}s$ (7.44)
$\quad\quad+C_{p}|\mathscr{G}|^{\frac{p}{2}}\int_{0}^{t}\mathbb{E}[(1+\sup_{0\leq t\leq T}|\widetilde{X}_{t}|^{p})]\mathrm{d}s$
$\leq C_{p}\left|\mathscr{G}\right|^{\frac{p}{2}}+C_{p}\int_{0}^{t}\mathbb{E}\left[\sup_{0\leq u\leq s}|\Delta X_{u}|\right]\mathrm{d}s.$
It follows from the Grönwall’s inequality that
$\mathbb{E}\left[\sup_{0\leq s\leq T}\left|\Delta X_{s}\right|^{p}\right]\leq C_{p}|\mathscr{G}|^{\frac{p}{2}}.$ (7.45)
Now we prove the second part of this theorem. If $|\mathscr{G}|=\frac{T}{n}$, then for $p\geq 2$, $t\geq 0$,
$\mathbb{P}\left(\sup_{0\leq t\leq T}|\Delta X_{t}|>t\right)=\mathbb{P}\left(\sup_{0\leq t\leq T}|\Delta X_{t}|^{p}>t^{p}\right)\leq t^{-p}\mathbb{E}\left[\sup_{0\leq t\leq T}|\Delta X_{t}|^{p}\right]\leq C_{p}n^{-\frac{p}{2}}t^{-p}.$ (7.46)
Let $t=n^{-\frac{1}{2}+\varepsilon}$,
$\mathbb{P}\left(\sup_{0\leq t\leq T}|\Delta X_{t}|>n^{-\frac{1}{2}+\varepsilon}\right)\leq C_{p}n^{-\varepsilon p}.$ (7.47)
By choosing $p=\frac{1}{\varepsilon}+2$, we have
$\sum_{n=1}^{\infty}\mathbb{P}\left(\sup_{0\leq t\leq T}|\Delta X_{t}|>n^{-\frac{1}{2}+\varepsilon}\right)\leq C_{p}\sum_{n=1}^{\infty}n^{-1-2\varepsilon}<+\infty.$ (7.48)
The desired result follows from the Borel-Cantelli lemma. ∎
7.2. Proofs of Propositions 5.1, 5.3 and 5.5
Proof of Proposition 5.1.
For the first part (1), note that
$J^{\mathscr{G}}-\tilde{J}=\left(\mathbb{E}\left[h(X^{\mathscr{G}}_{T})\right]-\mathbb{E}\left[h(\widetilde{X}_{T})\right]\right)+\sum_{i=0}^{n-1}\mathbb{E}\left[\int_{t_{i}}^{t_{i+1}}(r(s,X^{\mathscr{G}}_{s},a_{t_{i}})-\tilde{r}(s,\widetilde{X}_{s}))\mathrm{d}s\right]$ (7.49)
By Theorem 4.1, there is a constant $C_{p}$, such that
$\left|\mathbb{E}\left[h(X^{\mathscr{G}}_{T})\right]-\mathbb{E}\left[h(\widetilde{X}_{T})\right]\right|\leq C_{p}\|h\|_{C_{p}^{4}}|\mathscr{G}|.$ (7.50)
As for the second term on the right hand side of (7.49), note that if $t_{i}\leq s<t_{i+1}$, then
$\mathbb{E}\left[r(s,X^{\mathscr{G}}_{s},a_{t_{i}})-\tilde{r}(s,\widetilde{X}_{s})\right]$ $=\mathbb{E}\left[r(s,X^{\mathscr{G}}_{s},a_{t_{i}})-r(t_{i},X^{\mathscr{G}}_{t_{i}},a_{t_{i}})\right]+\mathbb{E}\left[r(t_{i},X^{\mathscr{G}}_{t_{i}},a_{t_{i}})-\tilde{r}(s,\widetilde{X}_{s})\right]$ (7.51)
$=\mathbb{E}\left[r(s,X^{\mathscr{G}}_{s},a_{t_{i}})-r(t_{i},X^{\mathscr{G}}_{t_{i}},a_{t_{i}})\right]+\mathbb{E}\left[\tilde{r}(t_{i},X^{\mathscr{G}}_{t_{i}})-\tilde{r}(t_{i},\widetilde{X}_{t_{i}})\right]$
$\quad\quad+\mathbb{E}\left[\tilde{r}(t_{i},\widetilde{X}_{t_{i}})-\tilde{r}(s,\widetilde{X}_{s})\right].$
Since $r\in C^{2,0}_{p}$, it follows from Itô’s lemma that
$|\mathbb{E}\left[r(s,X^{\mathscr{G}}_{s},a_{t_{i}})-r(t_{i},X^{\mathscr{G}}_{t_{i}},a_{t_{i}})\right]|\leq C_{p}\|r\|_{C_{p}^{2,0}}|\mathscr{G}|.$ (7.52)
Set $f=\tilde{r}(t_{i},\cdot)$ in Theorem 4.1, we have
$\left|\mathbb{E}\left[\tilde{r}(t_{i},X^{\mathscr{G}}_{t_{i}})-\tilde{r}(t_{i},\widetilde{X}_{t_{i}})\right]\right|\leq C_{p}\|\tilde{r}(t_{i},\cdot)\|_{C_{p}^{4}}|\mathscr{G}|.$ (7.53)
Using Itô’s formula, we have
$\left|\mathbb{E}\left[\tilde{r}(t_{i},\widetilde{X}_{t_{i}})-\tilde{r}(s,\widetilde{X}_{s})\right]\right|\leq C_{p}\|\tilde{r}\|_{C_{p}^{2}}|\mathscr{G}|.$ (7.54)
Combining these inequalities, we prove the desired result.
For the second part (2), we start at (7.49). By Theorem 4.2, there is a constant $C$, such that
$\left|\mathbb{E}\left[h(X^{\mathscr{G}}_{T})\right]-\mathbb{E}\left[h(\widetilde{X}_{T})\right]\right|\leq C\|h\|^{(l+2)}|\mathscr{G}|^{\mathcal{X}(l)}.$ (7.55)
As for the second term on the right hand side of (7.49), we still have (7.51). It follows from Lemma 7.1 that
$|\mathbb{E}\left[r(s,X^{\mathscr{G}}_{s},a_{t_{i}})-r(t_{i},X^{\mathscr{G}}_{t_{i}},a_{t_{i}})\right]|\leq C\|r\|_{T}^{(l)}|\mathscr{G}|^{\mathcal{X}(l)}.$ (7.56)
Set $f=\tilde{r}(t_{i},\cdot)$ in Theorem 4.2, we have
$\left|\mathbb{E}\left[\tilde{r}(t_{i},X^{\mathscr{G}}_{t_{i}})-\tilde{r}(t_{i},\widetilde{X}_{t_{i}})\right]\right|\leq C\|\tilde{r}(t_{i},\cdot)\|^{(l+2)}|\mathscr{G}|^{\mathcal{X}(l)}\leq C\sup_{0\leq t\leq T}\|\tilde{r}(t,\cdot)\|^{(l+2)}|\mathscr{G}|^{\mathcal{X}(l)}.$ (7.57)
Note that $\tilde{r}\in\mathcal{H}^{(l)}_{T}$, by the results of [23, Lemma 14.1.6], we have
$\left|\mathbb{E}\left[\tilde{r}(t_{i},\widetilde{X}_{t_{i}})-\tilde{r}(s,\widetilde{X}_{s})\right]\right|\leq C\|\tilde{r}\|_{T}^{(l)}|\mathscr{G}|^{\mathcal{X}(l)}.$ (7.58)
Combining these inequalities, we prove the desired result. ∎
Proof of Proposition 5.3.
For the variance part, note that
$\mathbb{E}\left[(S\circ V)^{2}\right]$ (7.59)
$=$ $\mathbb{E}\bigg{[}\bigg{(}\int_{0}^{T}S(\delta(t),X^{\mathscr{G}}_{\delta(t)},a_{\delta(t)})g(t,X^{\mathscr{G}}_{t},a_{\delta(t)})\mathrm{d}t+\int_{0}^{T}S(\delta(t),X^{\mathscr{G}}_{\delta(t)},a_{\delta(t)})(\partial_{x}V(t,X^{\mathscr{G}}_{t}))^{\top}\sigma(t,X^{\mathscr{G}}_{t},a_{\delta(t)})\mathrm{d}W_{t}\bigg{)}^{2}\bigg{]}$
$\leq$ $2\int_{[0,T]^{2}}\mathbb{E}\left[S(\delta(t),X^{\mathscr{G}}_{\delta(t)},a_{\delta(t)})g(t,X^{\mathscr{G}}_{t},a_{\delta(t)})S(\delta(s),X^{\mathscr{G}}_{\delta(s)},a_{\delta(s)})g(s,X^{\mathscr{G}}_{s},a_{\delta(s)})\right]\mathrm{d}s\mathrm{d}t$
$+2\int_{0}^{T}\mathbb{E}\left[\left|S(\delta(t),X^{\mathscr{G}}_{\delta(t)},a_{\delta(t)})(\partial_{x}V(t,X^{\mathscr{G}}_{t}))^{\top}\sigma(t,X^{\mathscr{G}}_{t},a_{\delta(t)})\right|^{2}\right]\mathrm{d}t.$
The desired result follows from Lemma 3.1.
For the bias part, it follows from Itô’s formula that
$\mathbb{E}^{W}\left[\int_{0}^{T}S(t,\widetilde{X}_{t})\left(\mathrm{d}V(t,\widetilde{X}_{t})+\tilde{r}(t,\widetilde{X}_{t})\mathrm{d}t\right)\right]$ (7.60)
$=\mathbb{E}^{W}\left[\int_{0}^{T}S(t,\widetilde{X}_{t})\left(\partial_{t}V+(\tilde{b}^{\bm{\pi}})^{\top}\partial_{x}V+\frac{1}{2}\operatorname*{Tr}\left((\tilde{\sigma}^{\bm{\pi}})^{2}\partial_{x}^{2}V\right)+\tilde{r}\right)(t,\widetilde{X}_{t})\mathrm{d}t\right]$
and
$\mathbb{E}[(S,V)]$ $=\sum_{i=0}^{n-1}\mathbb{E}\left[S(t_{i},X^{\mathscr{G}}_{t_{i}})\left(V(t_{i+1},X^{\mathscr{G}}_{t_{i+1}})-V(t_{i},X^{\mathscr{G}}_{t_{i}})+R^{\mathscr{G}}_{t_{i}}\right)\right]$ (7.61)
$=\mathbb{E}\bigg{[}\int_{0}^{T}S(\delta(t),X^{\mathscr{G}}_{\delta(t)})g(t,X^{\mathscr{G}}_{t},a_{\delta(t)})\mathrm{d}t\bigg{]}$
$=\mathbb{E}\left[\int_{0}^{T}S(t,X^{\mathscr{G}}_{t})g(t,X^{\mathscr{G}}_{t},a_{\delta(t)})\mathrm{d}t\right]-\mathbb{E}\left[\int_{0}^{T}\Delta S_{t}g(t,X^{\mathscr{G}}_{t},a_{\delta(t)})\mathrm{d}t\right],$
where $\Delta S_{t}:=S(t,X^{\mathscr{G}}_{t})-S(\delta(t),X^{\mathscr{G}}_{\delta(t)})$. We have
$\left|\mathbb{E}[S\circ V]-\mathbb{E}^{W}\left[\int_{0}^{T}S(t,\widetilde{X}_{t})\left(\mathrm{d}V(t,\widetilde{X}_{t})+\tilde{r}(t,\widetilde{X}_{t})\mathrm{d}t\right)\right]\right|$ (7.62)
$\leq\left|\mathbb{E}\left[\int_{0}^{T}Sg(t,X^{\mathscr{G}}_{t},a_{\delta(t)})\mathrm{d}t\right]-\mathbb{E}\left[\int_{0}^{T}\int_{A}Sg(t,\widetilde{X}_{t},a)\bm{\pi}(\mathrm{d}a|t,\widetilde{X}_{t})\mathrm{d}t\right]\right|+\left|\mathbb{E}\left[\int_{0}^{T}\Delta S_{t}g(t,X^{\mathscr{G}}_{t},a_{\delta(t)})\mathrm{d}t\right]\right|.$
The second term on the right hand side of above equation is bounded by $C_{p}\|S\|_{C_{p}^{2,0}}\|g\|_{C_{p}^{0,0}}|\mathscr{G}|$ and $C\|S\|_{T}^{(l)}\|g\|_{T}^{(0)}|\mathscr{G}|^{\mathcal{X}(l)}$ using the methods of Theorem 4.1 and Theorem 4.2, respectively. The desired results follow from the application of Proposition 5.1. ∎
Proof of Proposition 5.5.
For the variance part, note that
$\mathrm{Var}\left(S\Box V\right)$
$\leq\mathbb{E}\left[(S\Box V)^{2}\right]$
$=\mathbb{E}\left[\left(\sum_{i=1}^{n}S(t_{i-1},X^{\mathscr{G}}_{t_{i-1}},a_{t_{i}})\left(\int_{t_{i-1}}^{t_{i}}g_{1}(s,X^{\mathscr{G}}_{s},a_{t_{i-1}})\mathrm{d}s+\int_{t_{i-1}}^{t_{i}}(\partial_{x}V(s,X^{\mathscr{G}}_{s}))^{\top}\sigma(s,X^{\mathscr{G}}_{s},a_{t_{i-1}})\mathrm{d}W_{s}\right)^{2}\right)^{2}\right]$
$\leq C_{p}\|S\|_{C_{p}^{0,0}}^{2}\left(\|g_{1}\|_{C_{p}^{0,0}}^{2}+\|V\|_{C_{p}^{2}}^{2}\right)^{2}.$
For the bias part, it follows from Itô’s formula that
$S\Box V$ $=\sum_{i=1}^{n}S(t_{i-1},X^{\mathscr{G}}_{t_{i-1}},a_{t_{i}})\left(V(t_{i},X^{\mathscr{G}}_{t_{i}})-V(t_{i-1},X^{\mathscr{G}}_{t_{i-1}})+R^{\mathscr{G}}_{t_{i-1}}\right)^{2}$ (7.63)
$=\sum_{i=1}^{n}S(t_{i-1},X^{\mathscr{G}}_{t_{i-1}},a_{t_{i}})\left(\int_{t_{i-1}}^{t_{i}}g_{1}(s,X^{\mathscr{G}}_{s},a_{t_{i-1}})\mathrm{d}s+\int_{t_{i-1}}^{t_{i}}(\partial_{x}V(s,X^{\mathscr{G}}_{s}))^{\top}\sigma(s,X^{\mathscr{G}}_{s},a_{t_{i-1}})\mathrm{d}W_{s}\right)^{2}$
$=\sum_{i=1}^{n}S(t_{i-1},X^{\mathscr{G}}_{t_{i-1}},a_{t_{i}})\bigg{(}\bigg{(}\int_{t_{i-1}}^{t_{i}}g_{1}(s,X^{\mathscr{G}}_{s},a_{t_{i-1}})\mathrm{d}s\bigg{)}^{2}$
$\quad+2\int_{t_{i-1}}^{t_{i}}\big{(}g_{1}(s,X^{\mathscr{G}}_{s},a_{t_{i-1}})-g_{1}(t_{i-1},X^{\mathscr{G}}_{t_{i-1}},a_{t_{i-1}})\big{)}\mathrm{d}s\int_{t_{i-1}}^{t_{i}}(\partial_{x}V(s,X^{\mathscr{G}}_{s}))^{\top}\sigma(s,X^{\mathscr{G}}_{s},a_{t_{i-1}})\mathrm{d}W_{s}$
$\quad+2\Delta t_{i}g(t_{i-1},X^{\mathscr{G}}_{t_{i-1}},a_{t_{i-1}})\int_{t_{i-1}}^{t_{i}}(\partial_{x}V(s,X^{\mathscr{G}}_{s}))^{\top}\sigma(s,X^{\mathscr{G}}_{s},a_{t_{i-1}})\mathrm{d}W_{s}$
$\quad+\bigg{(}\int_{t_{i-1}}^{t_{i}}(\partial_{x}V(s,X^{\mathscr{G}}_{s}))^{\top}\sigma(s,X^{\mathscr{G}}_{s},a_{t_{i-1}})\mathrm{d}W_{s}\bigg{)}^{2}\bigg{)}.$
Note that the first and second terms are bounded by $C_{p}(\|S\|_{C_{p}^{2,0}}\|g_{1}\|_{C_{p}^{2,0}}^{2}+\|g_{1}\|_{C_{p}^{2,0}}\|V\|_{C_{p}^{2}})|\mathscr{G}|$ and $C(\|S\|_{T}^{(l)}(\|g_{1}\|_{T}^{(l)})^{2}+\|g_{1}\|_{T}^{(l)}\|V\|_{T}^{(l)})|\mathscr{G}|^{\mathcal{X}(l)}$ under different types of conditions, and third term is mean zero. For the last term, we have
$\mathbb{E}\left[\sum_{i=0}^{n-1}S(t_{i-1},X^{\mathscr{G}}_{t_{i-1}},a_{t_{i-1}})\left(\int_{t_{i-1}}^{t_{i}}(\partial_{x}V(s,X^{\mathscr{G}}_{s}))^{\top}\sigma(s,X^{\mathscr{G}}_{s},a_{t_{i-1}})\mathrm{d}W_{s}\right)^{2}\right]$ (7.64)
$=\mathbb{E}\left[\int_{0}^{T}S(\delta(t),X^{\mathscr{G}}_{\delta(t)},a_{\delta(t)})g_{2}(t,X^{\mathscr{G}}_{t},a_{\delta(t)})\mathrm{d}t\right].$
It converges to $\mathbb{E}\left[\int_{0}^{T}\widetilde{Sg}_{2}(t,\widetilde{X}_{t})\mathrm{d}t\right]$ with error bounds $C_{p}(\|S\|_{C_{p}^{2,0}}\|g_{2}\|_{C_{p}^{0,0}}+\|Sg_{2}\|_{C_{p}^{2,0}}+\|\widetilde{Sg_{2}}\|_{C_{p}^{2}}+\sup_{t\in[0,T]}\|\widetilde{Sg_{2}}(t,\cdot)\|_{C_{p}^{4}})|\mathscr{G}|$ and $C(\|S\|_{T}^{(l)}\|g_{2}\|_{T}^{(0)}+\|Sg_{2}\|_{T}^{(l)}+\|\widetilde{Sg_{2}}\|_{T}^{(l)}+\sup_{t\in[0,T]}\|\widetilde{Sg_{2}}(t,\cdot)\|^{(l)})|\mathscr{G}|^{\mathcal{X}(l)}$ under different types of conditions by the results of Proposition 5.1. In particular, if $S$ do not depend on $a$, then
$\mathbb{E}\left[\int_{0}^{T}\int_{A}S\partial_{x}V^{\top}\sigma\sigma^{\top}\partial_{x}V(t,\widetilde{X}_{t},a)\bm{\pi}(\mathrm{d}a|t,\widetilde{X}_{t})\mathrm{d}t\right]$ $=\mathbb{E}\left[\int_{0}^{T}S(t,\widetilde{X}_{t})\partial_{x}V(t,\widetilde{X}_{t})^{\top}(\tilde{\sigma}^{\bm{\pi}})^{2}(t,\widetilde{X}_{t})\partial_{x}V(t,\widetilde{X}_{t})\mathrm{d}t\right]$
$=\mathbb{E}\left[\int_{0}^{T}S(t,\widetilde{X}_{t})\mathrm{d}\left<V(\cdot,\widetilde{X}_{\cdot})\right>_{t}\right].$
This proves the desired results. ∎
7.3. Proofs of Examples 4.1 and 4.2
Proof of Example 4.1.
It suffices to work on $\tilde{\sigma}^{\bm{\pi}}$. Note that
$\partial_{x}^{s}(\tilde{\sigma}^{\bm{\pi}})^{2}(t,x)=\sum_{s_{1}+s_{2}=s}\int_{A}\partial_{x}^{s_{1}}\sigma\sigma^{\top}(t,x,a)\partial_{x}^{s_{2}}H(t,x,a)\mu(\mathrm{d}a)$ (7.65)
is bounded. The desired result follows from (7.70) and $\sigma\sigma^{\top}$ is uniformly elliptic. ∎
Proof of Example 4.2.
For simplicity, we only prove that $\tilde{\sigma}^{\bm{\pi}}$ has bounded derivatives with respect to $x$ for the case $d=m=1$. Note that
$\int_{\mathbb{R}}\sigma^{2}(t,x,a)\varphi(a|\mu(t,x),\Sigma(t,x))\mathrm{d}a=\int_{\mathbb{R}}\sigma^{2}(t,x,\mu(t,x)+\Sigma(t,x)u)\varphi(u)\mathrm{d}u:=\lambda(t,x),$ (7.66)
where $\varphi$ is the standard Gaussian density. Note that
$\partial_{x}\lambda^{\frac{1}{2}}$ $=\frac{1}{2}\lambda^{-\frac{1}{2}}\partial_{x}\lambda$ (7.67)
$\partial_{x}^{2}\lambda^{\frac{1}{2}}$ $=-\frac{1}{4}\lambda^{-\frac{3}{2}}(\partial_{x}\lambda)^{2}+\frac{1}{2}\lambda^{-\frac{1}{2}}\partial_{x}^{2}\lambda$
$\partial_{x}^{3}\lambda^{\frac{1}{2}}$ $=\frac{3}{8}\lambda^{-\frac{5}{2}}(\partial_{x}\lambda)^{3}-\frac{3}{4}\lambda^{-\frac{3}{2}}\partial_{x}\lambda\partial_{x}^{2}\lambda+\frac{1}{2}\lambda^{-\frac{1}{2}}\partial_{x}^{3}\lambda$
$\partial_{x}^{4}\lambda^{\frac{1}{2}}$ $=-\frac{15}{16}\lambda^{-\frac{7}{2}}(\partial_{x}\lambda)^{4}+\frac{9}{4}\lambda^{-\frac{5}{2}}(\partial_{x}\lambda)^{2}\partial_{x}^{2}\lambda-\frac{3}{4}\lambda^{-\frac{3}{2}}(\partial_{x}^{2}\lambda)^{2}-\lambda^{-\frac{3}{2}}\partial_{x}\lambda\partial_{x}^{3}\lambda+\frac{1}{2}\lambda^{-\frac{1}{2}}\partial_{x}^{4}\lambda$
Under the conditions of this proposition, we can exchange the order of integration and differentiation, therefore,
$\partial_{x}\lambda=\partial_{x}\int_{\mathbb{R}}\sigma^{2}\varphi(u)\mathrm{d}u=\int_{\mathbb{R}}2\sigma(\partial_{x}\sigma+\partial_{a}\sigma(\partial_{x}\mu+\partial_{x}\Sigma u))\varphi(u)\mathrm{d}u.$ (7.68)
It follows from Hölder’s inequality that
$|\partial_{x}\lambda|\leq 2\sqrt{\int_{\mathbb{R}}\sigma^{2}\varphi(u)\mathrm{d}u}\sqrt{\int_{\mathbb{R}}(\partial_{x}\sigma+\partial_{a}\sigma(\partial_{x}\mu+\partial_{x}\Sigma u))^{2}\varphi(u)\mathrm{d}u}\leq C\lambda^{\frac{1}{2}}.$ (7.69)
One can use the same method to obtain that for $s\leq 4$, there is a constant $C$ such that
$|\partial_{x}^{s}\lambda|\leq C(1+\lambda^{\frac{1}{2}}).$ (7.70)
Note that $\sigma\sigma^{\top}$ is uniformly elliptic. There is a $\varepsilon>0$ such that $\lambda^{-1}\leq\varepsilon^{-1}$. The desired result follows from (7.70) and (7.67). ∎
Appendix A Proofs of Lemmas 3.1 and 7.2
Proof of Lemma 3.1.
Throughout this proof, we denote by $C\geq 0$ a generic constant independent of $t_{i}$, which may take a different value at each occurrence. Note that $X^{\mathscr{G}}_{0}$ is deterministic and hence is $p$-th power integrable for all $p\geq 2$. Suppose that for some $i=0,\ldots,n-2$, (3.3) admits a unique strong solution $X^{\mathscr{G}}$ on $[0,t_{i}]$ satisfying ${\mathbb{E}}[\sup_{t\in[0,t_{i}]}|X^{\mathscr{G}}_{t}|^{p}]<\infty$ for all $p\geq 2$. Observe that by (H.1) and the Cauchy-Schwarz inequality, for all $p\geq 2$,
${\mathbb{E}}\left[\left(\int_{t_{i}}^{t_{i+1}}|b(t,0,a_{t_{i}})|{\mathrm{d}}t\right)^{p}\right]$ $\leq C(1+{\mathbb{E}}\left[d_{A}(a_{t_{i}},a_{0})^{p}\right])=C(1+{\mathbb{E}}\left[{\mathbb{E}}\left[d_{A}(a_{t_{i}},a_{0})^{p}\mid X^{\mathscr{G}}_{t_{i}}\right]\right])$
$=C\left(1+{\mathbb{E}}\left[\int_{A}d_{A}(a,a_{0})^{p}\bm{\pi}(da|t_{i},X^{\mathscr{G}}_{t_{i}})\right]\right)\leq C\left(1+{\mathbb{E}}\left[|X^{\mathscr{G}}_{t_{i}}|^{p}\right]\right)<\infty.$
Similar arguments show that
${\mathbb{E}}\left[\left(\int_{t_{i}}^{t_{i+1}}|\sigma(t,0,a_{t_{i}})|^{2}{\mathrm{d}}t\right)^{p/2}\right]\leq{\mathbb{E}}\left[\int_{t_{i}}^{t_{i+1}}|\sigma(t,0,a_{t_{i}})|^{p}{\mathrm{d}}t\right]<\infty.$
Since $x\mapsto b(t,x,a)$ and $x\mapsto\sigma(t,x,a)$ are Lipschitz continuous uniformly in $(t,a)$, by [40, Theorems 3.3.1 and 3.4.3], (3.3) with the initial state $X^{\mathscr{G}}_{t_{i}}$ at $t_{i}$ admits a unique strong solution $X^{\mathscr{G}}$ on $[t_{i},t_{i+1}]$ satisfying ${\mathbb{E}}[\sup_{t\in[t_{i},t_{i+1}]}|X^{\mathscr{G}}_{t}|^{p}]<\infty$ for all $p\geq 2$. This along with an inductive argument shows the desired well-posedness result.
Apply the notation $\delta(u):=t_{i}$ for $u\in[t_{i},t_{i+1})$, then it satisfies that
$X^{\mathscr{G}}_{s}=x_{0}+\int_{t}^{s}b(u,X^{\mathscr{G}}_{u},a_{\delta(u)})\mathrm{d}u+\int_{t}^{s}\sigma(u,X^{\mathscr{G}}_{u},a_{\delta(u)})\mathrm{d}W_{u}.$ (A.1)
Given that $b(u,X^{\mathscr{G}}_{u},a_{\delta(u)})$ and $\sigma(u,X^{\mathscr{G}}_{u},a_{\delta(u)})$ have finite moments, there is a constant $C_{p}>0$, depending only on $p$, such that
$\mathbb{E}\left[\left|X^{\mathscr{G}}_{s}\right|^{p}\right]$ $\leq C_{p}\left|x_{0}\right|^{p}+C_{p}\int_{t}^{s}\mathbb{E}\left[\left|b(u,X^{\mathscr{G}}_{u},a_{\delta(u)})\right|^{p}\right]\mathrm{d}u$ (A.2)
$+C_{p}\int_{t}^{s}\mathbb{E}\left[\left|\sigma(u,X^{\mathscr{G}}_{u},a_{\delta(u)})\right|^{p}\right]\mathrm{d}u.$
Noting that by (H.1), for $h=b$ or $\sigma$, it satisfies that
$\mathbb{E}\left[\left|h(u,X^{\mathscr{G}}_{u},a_{\delta(u)})\right|^{p}\right]$ $\leq\mathbb{E}\left[C_{p}\left(1+\left|X^{\mathscr{G}}_{u}\right|^{p}+d_{A}(a_{0},a_{\delta(u)})^{p}\right)\right]$ (A.3)
$\leq C_{p}\left(1+\mathbb{E}\left[\left|X^{\mathscr{G}}_{u}\right|^{p}\right]+\mathbb{E}\left[\left|X^{\mathscr{G}}_{\delta(u)}\right|^{p}\right]\right),$
where the second inequality follows from taking conditional expectation $\mathbb{E}[\cdot|X^{\mathscr{G}}_{\delta(u)}]$ and using Condition (H.13). Substituting (A.3) into (A.2), we obtain
$\mathbb{E}\left[\left|X^{\mathscr{G}}_{s}\right|^{p}\right]\leq C_{p}\left|x_{0}\right|^{p}+C_{p}\int_{t}^{s}\left(1+\mathbb{E}\left[\left|X^{\mathscr{G}}_{u}\right|^{p}\right]+\mathbb{E}\left[\left|X^{\mathscr{G}}_{\delta(u)}\right|^{p}\right]\right)\mathrm{d}u.$ (A.4)
Choosing $s=\delta(s)$ in (A.4), and adding it with (A.4), we obtain
$\mathbb{E}\left[\left|X^{\mathscr{G}}_{s}\right|^{p}\right]$ $+\mathbb{E}\left[\left|X^{\mathscr{G}}_{\delta(s)}\right|^{p}\right]$
$\leq$ $2C_{p}\left|x_{0}\right|^{p}+2C_{p}\int_{t}^{s}\left(1+\mathbb{E}\left[\left|X^{\mathscr{G}}_{u}\right|^{p}\right]+\mathbb{E}\left[\left|X^{\mathscr{G}}_{\delta(u)}\right|^{p}\right]\right)\mathrm{d}u.$
The desired result follows from the Grönwall’s inequality. ∎
Proof of Lemma 7.2.
We first prove $\mathbb{E}^{W}[\sup_{t\in[0,T]}|X^{\mathscr{G}}_{t}|^{p}]<\infty$ for all $p\geq 2$. Observe that by (H.5), for all $p\geq 2$,
$\mathbb{E}^{W}\left[\left(\int_{t_{i}}^{t_{i+1}}|b(t,0,a_{t_{i}})|\mathrm{d}t\right)^{p}\right]\leq C<\infty.$ (A.5)
Similarly, we have
$\mathbb{E}^{W}\left[\left(\int_{t_{i}}^{t_{i+1}}|\sigma(t,0,a_{t_{i}})|^{2}\mathrm{d}t\right)^{p/2}\right]\leq\mathbb{E}^{W}\left[\int_{t_{i}}^{t_{i+1}}|\sigma(t,0,a_{t_{i}})|^{p}\mathrm{d}t\right]\leq C<\infty.$ (A.6)
Note that $\mathbb{E}^{W}\left[\cdot\right]=\mathbb{E}[\cdot|\widetilde{\mathcal{F}}_{0}]$, and all Itô integrals appear here equal to 0 when taking the conditional expectation $\mathbb{E}[\cdot|\widetilde{\mathcal{F}}_{0}]$. Analogous to the proofs of Lemma 3.1 and [40, Theorem 3.4.3], we obtain that $\mathbb{E}^{W}[\sup_{t\in[0,T]}|X^{\mathscr{G}}_{t}|^{p}]<\infty$ for all $p\geq 2$.
Next, we aim to prove that the moments of $X^{\mathscr{G}}_{t}$ is independent of $|\mathscr{G}|$ and obtain (7.10). Use the notation $\delta(u):=t_{i}$ for $u\in[t_{i},t_{i+1})$, then it satisfies that
$X^{\mathscr{G}}_{s}=x_{0}+\int_{0}^{s}b(u,X^{\mathscr{G}}_{u},a_{\delta(u)})\mathrm{d}u+\int_{0}^{s}\sigma(u,X^{\mathscr{G}}_{u},a_{\delta(u)})\mathrm{d}W_{u}.$ (A.7)
Note that $b(u,X^{\mathscr{G}}_{u},a_{\delta(u)})$ and $\sigma(u,X^{\mathscr{G}}_{u},a_{\delta(u)})$ have finite moments by (H.5). There is a constant $C_{p}>0$, depending only on $p$, such that
$\mathbb{E}^{W}\left[\left|X^{\mathscr{G}}_{s}\right|^{p}\right]$ $\leq C_{p}\left|x_{0}\right|^{p}+C_{p}\int_{0}^{s}\mathbb{E}^{W}\left[\left|b(u,X^{\mathscr{G}}_{u},a_{\delta(u)})\right|^{p}\right]\mathrm{d}u$ (A.8)
$+C_{p}\int_{0}^{s}\mathbb{E}^{W}\left[\left|\sigma(u,X^{\mathscr{G}}_{u},a_{\delta(u)})\right|^{p}\right]\mathrm{d}u.$
Noting that by condition (H.5), for $h=b$ or $\sigma$, it satisfies that
$\mathbb{E}^{W}\left[\left|h(u,X^{\mathscr{G}}_{u},a_{\delta(u)})\right|^{p}\right]$ $\leq\mathbb{E}^{W}\left[C_{p}\left(1+\left|X^{\mathscr{G}}_{u}\right|^{p}\right)\right].$ (A.9)
Substituting (A.9) into (A.8), we obtain
$\mathbb{E}^{W}\left[\left|X^{\mathscr{G}}_{s}\right|^{p}\right]\leq C_{p}\left|x_{0}\right|^{p}+C_{p}\int_{t}^{s}\left(1+\mathbb{E}^{W}\left[\left|X^{\mathscr{G}}_{u}\right|^{p}\right]\right)\mathrm{d}u.$ (A.10)
The desired result follows from the Grönwall’s inequality. ∎